### v3

# finite free probability and $S$ transforms of Jacobi processes

Abstract.
In this paper, we study the $S$ transforms of Jacobi processes in the frameworks of free and finite free probability theories. We begin by deriving a partial differential equation satisfied by the free $S$ transform of the free Jacobi process, and we provide a detailed analysis of its characteristic curves. We turn next our attention to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamic of its roots, referred to as the frozen Jacobi process. In particular, we prove, for a specific set of parameters, that the former aligns up to a Szegö variable transformation with the Hermite unitary polynomial. We also provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials. Finally, we establish the convergence of the frozen Jacobi process to the free Jacobi process in high dimensions by using the finite free S transform. In doing so, we prove a general result, interesting in its own, on the convergence of the finite differences of the finite free $S$ transform, which paves the way to obtain asymptotics of differential-difference equations satisfied by time-dependent finite free S transforms of polynomial sequences with positive roots.
Key words and phrases:
Hermitian and free Jacobi processes; Free and finite free S transforms; Averaged characteristic polynomial
1. Introduction
1.1. Finite free probability (FFP)
Introduced and studied in [MR4408504], FFP has gained a lot of attention in recent years, see [arizmendi2018cumulants, arizmendi2024s, MR4408504, MR4925017]. It provides an original perspective on certain operations on polynomials that were defined long before the seminal work [MR4408504]. Actually, FFP involves convolution operations on averaged characteristic polynomials of complex random matrices, whose laws have prescribed symmetry invariances, as well as the computations of their root distributions. Important questions deal with asymptotics of these root distributions when the degrees of their corresponding polynomials tend to infinity. In this regime, the FFP convolutions tend to the convolutions already introduced by Voiculescu in his theory of free probability. Further investigations concern finite free counterparts of results and transforms of prime importance in Voiculescu’s theory (see e.g. [MR4925017], [MR4912666], [arizmendi2024s], [JKM25], [Kab25]).
In [MR4408504], finite versions of the free additive, the free rectangular and the free multiplicative convolutions were studied, though the latter is attributed to [MR4285332]. The finite free unitary convolution was also defined in [MR4285332], and the corresponding central limit theorem was proved there. In this respect, the limiting polynomial is given by the so-called unitary Hermite polynomial, which was shown in [MR4912666] to coincide (up to rescaling) with the characteristic polynomial of the Brownian motion on the unitary group. In that paper, it was also proved that the corresponding empirical root distribution converges weakly to the spectral distribution of the free unitary Brownian motion [MR1430721].
1.2. Crystallization and the $\beta$-Jacobi process
Among finite versions of various results in random matrix and free probability theories figure the analytic extrapolation to any $\beta>0$ of finite free convolutions and the $\beta$-corners processes [MR4073197]. In particular, this extrapolation shows that the elementary symmetric polynomials in the roots of averaged characteristic polynomials of random matrices with prescribed law-invariance properties do not depend on $\beta$. This fact was already noted in earlier papers for classical Hermite/Laguerre/Jacobi ensembles and opens the way to taking the limit $\beta\rightarrow+\infty$. Doing so leads to the extension of the so-called crystallization beyond the eigenvalues of classical matrix-valued processes.
The $\beta$-Jacobi process, introduced and studied in [MR2719370], is an extrapolation to all $\beta>0$ of the eigenvalues of the real symmetric and the complex Hermitian Jacobi processes. For any $\beta>0$, its distribution converges weakly as time $t\rightarrow+\infty$ to the corresponding $\beta$-Jacobi ensemble. Besides, when $\beta=2$, the underlying Hermitian matrix model converges in the large-size limit to the free Jacobi process [MR2384475]. The moments of this matrix model and their large size limits were computed in a series of papers [MR3842169, MR4173022, MR4954166]. However, the Lebesgue decomposition of its spectral distribution was determined in [MR3812867]. Additionally, the expressions for the $R$ and $S$ transforms of the free Jacobi process associated with a specific parameter set were derived in [MR3679612] using the Lagrange inversion formula. In this respect, note that partial differential equations for $R$ and $S$ transforms relative to hydrodynamic limits of Dyson and Wishart processes were derived in [MR4446587] and solved using the method of characteristics.
As to the crystallization of the $\beta$-Jacobi process, it still holds as shown in [MR4673381]. Yet, it does not follow from the approach undertaken in [MR4073197] since the marginal distribution of this random particle system does not satisfy the required invariance properties. Rather, a combination of various results proved in [MR4092353], [MR4673381] and [MR4781073] shows that the averaged elementary symmetric polynomials in the rescaled process at any fixed time $t$ do not depend on $\beta$. Equivalently, the averaged characteristic polynomial coincides with that of the deterministic Jacobi particle system, referred to as the frozen Jacobi process. From a dynamical perspective, the differential system satisfied by the latter is equivalent to the inverse Jacobi-heat equation for the former [MR4781073]. Apparently, such equivalence was noticed for the first time by T. Tao [^1] https://terrytao.wordpress.com/2017/10/17/heat-flow-and-zeroes-of-polynomials/ for polynomial roots undergoing the heat flow and was motivated by an application to analytic number theory. A few years later, other root dynamics were derived in the unpublished paper [FFPP], which relied on determinantal formulas describing various finite free convolutions. Actually, these dynamics are governed by the drifts of eigenvalue processes of the Dyson, Laguerre and Hermitian Jacobi processes. However, the underlying matrix model from which the last dynamics arise is not the Hermitian Jacobi process as explained in [FFPP] (see section 5.2). For the sake of completeness, let us close this paragraph by recalling the following related results:
  • The long-time behaviour of the frozen Dyson, Laguerre and Jacobi processes is governed by the zeroes of Hermite, Laguerre and Jacobi polynomials (see e.g. [MR4781073]).
  • The dynamics of the roots of the Hermite unitary polynomial (which lie on the unit circle) were determined in [MR4285332]. In particular, their angles form the frozen process corresponding to the unitary $\beta$-Dyson process.
1.3. (Finite) free S transforms
To clarify connections between finite free probability and Voiculescu’s theory, finite analogues of various transforms central to free probability have been defined, including the $S$-transform [bercovici1993free]. The first proposal for the finite free $S$- transform is due to Marcus [MR4408504], followed a few years later by another proposal, motivated by the CLT for finite free multiplicative convolution [arizmendi2024s]. Marcus’s definition is general enough to cover the finite free multiplicative convolutions of distributions supported on the positive half-line and on the circle, but is much less practical for computations. For this reason, we prefer to work with the definition proposed in [arizmendi2024s] in the present work.
Recall that if $\mu$ is a probability measure on $\mathbb{R}_{+}$ different from the dirac measure at 0 and $\psi_{\mu}\colon\mathbb{C}_{+}\to\mathbb{C}$ is the moment transform of $\mu$ defined over the complex plane minus the positive half-line, then $\psi_{\mu}$ is univalent on the left upper-half plane $i\mathbb{C}_{+}$. Besides $\psi_{\mu}(i\mathbb{C}_{+})$ is a domain contained in the disc with diameter $(-1+\mu({0}),0)$ and $(-1+\mu({0}),0)\subset\psi_{\mu}(i\mathbb{C}_{+})$. The free $S$-transform of $\mu$ by, for $z$ in the vicinity of $(-1+\mu({0}),0)$ is then defined by
$S_{\mu}(z)=\frac{1+z}{z}\psi_{\mu}^{(-1)}(z).$
Moreover, it factorizes under the free multiplicative convolution of distributions on the positive half-line.
1.4. Our contributions
In the present work, we prove out-of-equilibrium equivalents of the previously stated results. More precisely, we investigate the $S$-transforms of Jacobi processes within the frameworks of both free and finite free probability theories. We begin by deriving a partial differential equation (PDE) that the $S$-transform of the free Jacobi process satisfies, as stated in Theorem 3.1.1. Once we do, we provide a detailed analysis of its stationary analytic solutions around the origin in relation to the angle operator between two free orthogonal projections, viewed as an operator in the compressed probability space. For the sake of simplicity and motivated by [MR3679612], we determine the characteristic curves of this PDE for the special parameter values $\lambda=1$ and $\theta=1/2$ and obtain an explicit expression of the S transform locally around the origin.
Our focus is then turned to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamics of its roots. In this respect, we write a direct proof of the fact that the former satisfies the inverse Jacobi-heat equation and prove further that the latter are contained in the interval $(0,1)$. In particular, when $\lambda=1$ and $\theta=1/2$, we show that the averaged characteristic polynomial at time $t$ aligns, up to a Szegö type transformation, with the Hermite unitary polynomial at time $2t$. This result is indeed the finite free analogue of Corollary 3.3 from [MR3071702]. Additionally, we provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials, valid for any initial data. To this end, we appeal to the dual Cauchy identity for multivariate Jacobi polynomials and to the semi-group density of the Jacobi eigenvalue process. Doing so allows us to use the mutual orthogonality of these polynomials, leading to the sought expansion.
Finally, we establish that the crystallization dynamics of the Jacobi process tend, in the regime of high degree and high dimension, towards the free Jacobi process and write two proofs of this result. The first one relies on the equi-continuity of the moment sequence of the empirical root distribution and on the fact that the sequence of the moments of the free Jacobi process is uniquely determined by the differential system it satisfies. While the second proof is based on the finite free equivalent of the $T$-transform in Voiculescu’s theory. This alternative approach is noteworthy because it requires a general technical result that we prove below regarding the convergence of the finite $T$-transform and of its discrete derivative. Once proved, we derive a differential-difference equation for the $T$-transform of the Hermitian Jacobi process and recover in the large size regime, the previously derived PDE for the free S transform.
2. Reminder: Jacobi Processes
Since our concern will be on the Hermitian Jacobi process and on its large-size limit, the free Jacobi process, we recall some facts needed in the sequel.
2.1. The Hermitian Jacobi process
Let $(Y_{t}(d))_{t\geq 0}$ be a unitary Brownian motion of size $d$ and variance $t$. Let $P_{m}$ and $Q_{p}$ be two diagonal orthogonal projections of rank $m$ and $p$ respectively, with $m\leq p$. The Hermitian Jacobi process $(X_{t}(m))_{t\geq 0}$ of parameters $(r,s,m)$ is the $m\times m$ matrix-valued stochastic process defined by:
$X_{t}(m)\oplus 0_{d-m}=P_{m}Y_{t}(d)Q_{p}Y_{t}^{\star}(d)P_{m}$
where
$r:=p-m,\quad s=d-p-m:=q-m.$
Its eigenvalues process $(\lambda_{t}^{i},t\geq 0)_{1\leq i\leq m}$ of $(X_{t})_{t\geq 0}$ satisfy the following stochastic differential system:
$d\lambda_{t}^{i}=\sqrt{2\lambda_{t}^{i}(1-\lambda_{t}^{i})}dB_{t}^{i}+\left[p-(p+q)\lambda_{t}^{i}+\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})+\lambda_{t}^{j}(1-\lambda_{t}^{i})}{\lambda_{t}^{i}-\lambda_{t}^{j}}\right]dt,$ (1)
where $(B_{t}^{i},t\geq 0)_{1\leq i\leq m}$ are independent Brownian motions [MR3842169]. Up to the time change $t\mapsto 2t$, this random particle system is an instance of the $\beta$-Jacobi process studied in [MR2384475] corresponding to the value $\beta=2$. In particular, Corollary 2.1 shows that the eigenvalues process is globally defined in time provided that $p\wedge q>m-1/2$. Corollary 7 in [MR3296535] shows that the eigenvalues of $X_{t}$ are almost surely distinct and that global existence holds even under the weaker condition $p\wedge q>m-1$. Consequently, Theorem 4.1 in [MR2384475] extends to $p\wedge q>m-1$ so that the probability distribution of the eigenvalues process $(\lambda_{t}^{i},t\geq 0)_{1\leq i\leq m}$ has a density with respect to Lebesgue measure in $\mathbb{R}^{m}$, for any $p\wedge q>m-1$.
2.2. The free Jacobi process
The free Jacobi process was introduced in [MR2384475] as the large-size limit of the Hermitian Jacobi process $(X_{t/d(m)}(m))_{t\geq 0}$ in the asymptotic regime:
$\lambda\theta=\lim_{m\rightarrow+\infty}\frac{m}{d(m)}\in(0,1],\quad\lambda=\lim_{m\rightarrow+\infty}\frac{m}{p(m)}>0.$ (2)
For any time $t\geq 0$, we can define it in an abstract way as
$X_{t}:=PY_{t}QY_{t}^{\star}P$
where $P,Q$ are free (in Voiculescu’s sense) orthogonal projections in a non commutative probability space $(\mathcal{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, and $(Y_{t})_{t\geq 0}$ is the so-called free unitary Brownian motion [MR1430721]. Viewed as an operator in the compressed probability space $(P\mathcal{A}P,\tau/\tau(P))$, the Cauchy transform $G_{t}^{(\lambda,\theta)}$ of its spectral distribution satisfies the following non-linear partial differential equation (PDE):
$\partial_{t}G_{t}^{(\lambda,\theta)}=\partial_{z}\left\{[(1-2\lambda\theta)z-\theta(1-\lambda)]G_{t}^{(\lambda,\theta)}+\lambda\theta z(z-1)[G_{t}^{(\lambda,\theta)}]^{2}\right\}.$
Consequently, the moment generating function
$\mathcal{M}_{t}^{(\lambda,\theta)}(z)=\frac{1}{z}G_{t}^{(\lambda,\theta)}\left(\frac{1}{z}\right)$ (3)
satisfies:
$\partial_{t}\mathcal{M}_{t}^{(\lambda,\theta)}=-z\partial_{z}\left\{[(1-2\lambda\theta)-\theta(1-\lambda)z]\mathcal{M}_{t}^{(\lambda,\theta)}+\lambda\theta(1-z)[\mathcal{M}_{t}^{(\lambda,\theta)}]^{2}\right\},$
which reduces when $\lambda=1,\theta=1/2,$ to
$\partial_{t}\mathcal{M}_{t}^{(1,1/2)}=-\frac{z}{2}\partial_{z}\left\{(1-z)[\mathcal{M}_{t}^{(1,1/2)}]^{2}\right\}.$
The solution of this equation subject to the initial data
$\mathcal{M}_{0}^{(1,1/2)}(z)=\frac{1}{1-z}$
was determined in [MR3071702] and used in [MR3679612] to derive expansions of the corresponding $R$ and $S$ transforms, based on Lagrange inversion formula.
Other degenerate limiting regimes were investigated in [MR4751304] for the Jacobi eigenvalue process, and gave rise to Marchenko-Pastur and Wigner distributions. At the root level, these regimes correspond to the ones studied in [MR1342385].
3. $S$ transform of the free Jacobi Process
This section is devoted to the study of the S transform of the free Jacobi process. More precisely, Theorem 3.1.1 below provides a PDE satisfied by this analytic (around the origin) map, whose stationary solutions are shown to coincide with the S transform of the angle operator $PQP$ between two free orthogonal projections (viewed as an operator in the compressed probability space). Though this coincidence is expected because the free Jacobi process converges weakly to $PQP$, it is not straightforward from a partial differential equation perspective. For that reason, we find it better to check it using standard (by now) facts from free probability theory, such as the Nica-Speicher convolution semi-group and the relation between the S and the free cumulant generating functions.
3.1. A PDE for the $S$ transform
For any time $t\geq 0$, denote $\eta_{t}^{(\lambda,\theta)}$ the local inverse around zero of
$\psi_{t}^{(\lambda,\theta)}:=\mathcal{M}_{t}^{(\lambda,\theta)}-1,$
where we recall that $\mathcal{M}_{t}^{(\lambda,\theta)}$ is the moment generating function of the free Jacobi process at time $t$. Since this nonnegative operator is non-degenerate (its first moment is non-zero), its $S$ transform $\mathcal{S}_{t}^{(\lambda,\theta)}$ is defined locally by
$\mathcal{S}_{t}^{(\lambda,\theta)}(z):=\frac{1+z}{z}\eta_{t}^{(\lambda,\theta)}(z),$
where the value at $z=0$ is obtained after removing the singularity at $z=0$. Then the PDE (3) entails:
Theorem 3.1.1.
The $S$ transform of the free Jacobi process satisfies locally around the origin the PDE:
$\partial_{t}\mathcal{S}_{t}^{(\lambda,\theta)}(z)=(2\lambda\theta z+1)\mathcal{S}_{t}^{(\lambda,\theta)}-\theta(1+2\lambda z)[\mathcal{S}_{t}^{(\lambda,\theta)}(z)]^{2}-\frac{\theta z(1+\lambda z)}{2}\partial_{z}[\mathcal{S}_{t}^{(\lambda,\theta)}(z)]^{2}.$ (4)
Proof.
Since $\psi_{t}^{(\lambda,\theta)}[\eta_{t}^{(\lambda,\theta)}(z)]=z$, then the chain rule entails:
$1$ $=$ $\partial_{z}\psi_{t}^{(\lambda,\theta)}[\eta_{t}^{(\lambda,\theta)}(z)]\partial_{z}\eta_{t}^{(\lambda,\theta)}(z)$
$0$ $=$ $\partial_{t}\psi_{t}^{(\lambda,\theta)}[\eta_{t}^{(\lambda,\theta)}(z)]+\partial_{t}[\eta_{t}^{(\lambda,\theta)}](z)\partial_{z}\psi_{t}^{(\lambda,\theta)}[\eta_{t}^{(\lambda,\theta)}(z)].$
It follows that
$\partial_{t}[\eta_{t}^{(\lambda,\theta)}]=-\partial_{t}\psi_{t}^{(\lambda,\theta)}[\eta_{t}^{(\lambda,\theta)}]\partial_{z}[\eta_{t}^{(\lambda,\theta)}]$
$=[\eta_{t}^{(\lambda,\theta)}]\partial_{z}\left\{[(1-2\lambda\theta)-\theta(1-\lambda)z](\psi_{t}^{(\lambda,\theta)}+1)+\lambda\theta(1-z)[(\psi_{t}^{(\lambda,\theta)}+1)]^{2}\right\}[\eta_{t}^{(\lambda,\theta)}]$
$\times\partial_{z}[\eta_{t}^{(\lambda,\theta)}]$
$=[\eta_{t}^{(\lambda,\theta)}]\partial_{z}\left\{(1+z)\left[(1-2\lambda\theta)-\theta(1-\lambda)\eta_{t}^{(\lambda,\theta)}\right]+\lambda\theta(z+1)^{2}\left(1-\eta_{t}^{(\lambda,\theta)}\right)\right\}$
$=[\eta_{t}^{(\lambda,\theta)}]\partial_{z}\left\{(1-\lambda\theta+z+\lambda\theta z^{2})-\theta(1+(1+\lambda)z+\lambda z^{2})\eta_{t}^{(\lambda,\theta)}\right\}$
$=(2\lambda\theta z+1)[\eta_{t}^{(\lambda,\theta)}]-\theta(1+\lambda+2\lambda z)\left[\eta_{t}^{(\lambda,\theta)}\right]^{2}-\frac{\theta}{2}(1+z)(1+\lambda z)\partial_{z}\left[\eta_{t}^{(\lambda,\theta)}\right]^{2}.$
Multiplying both sides by $(1+z)/z$ and using the fact that
$\partial_{z}\left[\mathcal{S}_{t}^{(\lambda,\theta)}\right]^{2}$ $=\frac{(1+z)^{2}}{z^{2}}\partial_{z}\left(\eta_{t}^{(\lambda,\theta)}\right)^{2}-2\frac{1+z}{z^{3}}\left(\eta_{t}^{(\lambda,\theta)}\right)^{2}$
$=\frac{(1+z)^{2}}{z^{2}}\partial_{z}\left(\eta_{t}^{(\lambda,\theta)}\right)^{2}-\frac{2}{z(1+z)}\left(\mathcal{S}_{t}^{(\lambda,\theta)}\right)^{2},$
we get the desired result. ∎
3.2. Stationary analytic (around the origin) solutions of (4)
These are the solutions of (4) which do not depend on time. As such, they satisfy the ordinary differential equation:
$(2\lambda\theta z+1)\mathcal{S}^{(\lambda,\theta)}-\theta(1+2\lambda z)[\mathcal{S}^{(\lambda,\theta)}(z)]^{2}-\frac{\theta z(1+\lambda z)}{2}\partial_{z}[\mathcal{S}^{(\lambda,\theta)}]^{2}(z)=0.$
The zero function is obviously the trivial solution. Otherwise, any non-zero stationary analytic solution satisfies
$\theta z(1+\lambda z)\partial_{z}\mathcal{S}^{(\lambda,\theta)}(z)+\theta(1+2\lambda z)\mathcal{S}^{(\lambda,\theta)}(z)=1+2\lambda\theta z.$
Noting that the LHS of this ODE may be written as $\partial_{z}[\theta z(1+\lambda z)\mathcal{S}^{(\lambda,\theta)}(z)]$, we readily see that
$\mathcal{S}^{(\lambda,\theta)}(z)=\frac{\lambda\theta z^{2}+z+\theta c}{\theta z(\lambda z+1)},\quad z\in\mathbb{C}\setminus\{0,-\frac{1}{\lambda}\}\ {\rm and}\ c\in\mathbb{C}.$
But analyticity around the origin forces $c=0$ and we end up (after removing the singularity at $z=0$) with
$\mathcal{S}^{(\lambda,\theta)}(z)=\frac{\lambda\theta z^{2}+z}{\theta z(\lambda z+1)}=\frac{\lambda\theta z+1}{\lambda\theta z+\theta}.$
This is the $S$ transform $S_{Q}$ of an orthogonal projection $Q$ of rank $\theta$ taken at $\lambda\theta z$. Equivalently, it is the $S$ transform of the weak limit as $t\rightarrow+\infty$ of the spectral distribution of the free Jacobi process. The latter may be defined as the spectral distribution of the angle operator $PQP$ between two free orthogonal projections $P$ and $Q$ in a non commutative probability space $(\mathscr{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, viewed as an operator in the compressed probability space $(P\mathscr{A}P,\tau/\tau(P))$. Using the Nica-Speicher convolution semi-group, see [MR2266879], Lecture 14, the Voiculescu’s $R$ transform $\mathcal{R}_{PQP}$ of $PQP$ in the compressed space is given by
$R_{Q}(\lambda\theta z)$
where $R_{Q}$ is the Voiculescu’s $R$ transform of $Q$ in $(\mathscr{A},\tau)$, (see [MR2266879], page 211). Let
$\mathcal{C}_{PQP}(z):=z\mathcal{R}_{PQP}(z),\quad C_{Q}(z):=zR_{Q}(z),$
be the shifted Voiculescu R transforms (also known as free cumulant generating functions) of $PQP$ and of $Q$, respectively. Then ([MR2266879], Lecture 16)
$\mathcal{C}_{PQP}(z)=zR_{Q}(\lambda\theta z)=\frac{1}{\lambda\theta}C_{Q}(\lambda\theta z).$
Consequently, the local inverse around $z=0$ of $\mathcal{C}_{PQP}$ is given by
$\mathcal{C}_{PQP}^{-1}(z)=\frac{1}{\lambda\theta}C_{Q}^{-1}(\lambda\theta z).$
Finally, the functional relations
$\mathcal{C}_{PQP}^{-1}(z)=z\mathcal{S}_{PQP}(z),\quad C_{Q}^{-1}(z)=zS_{Q}(z),$
yield
$z\mathcal{S}_{PQP}(z)=\frac{\lambda\theta z}{\lambda\theta}S_{Q}(\lambda\theta z)\quad\Leftrightarrow\quad\mathcal{S}_{PQP}(z)=S_{Q}(\lambda\theta z)=\mathcal{S}^{(\lambda,\theta)}(z).$
In a nutshell,
$\mathcal{S}^{(\lambda,\theta)}(z)=\lim_{t\rightarrow+\infty}\mathcal{S}_{t}^{(\lambda,\theta)}(z):=\mathcal{S}_{\infty}^{(\lambda,\theta)}(z)$
is the non-trivial analytic stationary solution of the PDE (4).
3.3. Analysis of the Characteristic curves for $\lambda=1,\theta=1/2$
In this paragraph, we restrict our attention to the special parameter set $\theta=1/2$, $\lambda=1$, and assume for sake of simplicity that $\mathcal{S}^{(1,1/2)}_{0}(z)=1$ (in which case the initial spectral distribution of the free Jacobi process is $\delta_{1}$). This restriction is mainly motivated by the fact that the moment-generating function $\mathcal{M}_{t}^{(1,1/2)}$ admits an explicit expression, as proved in [MR3071702], from which a series expansion of $\mathcal{S}_{t}^{(1,1/2)}$ was derived in [MR3679612] making use of Lagrange inversion formula. As we shall see below, the analysis of the Characteristic curves of the pde (4) in the special case corresponding to $\lambda=1,\theta=1/2$ leads to a considerably more tractable expression of $\mathcal{S}^{(1,1/2)}_{t}$ than the one displayed in Proposition 5.2. from [MR3679612].
Proposition 3.3.1.
Let $t\geq 0$. The S transform $\mathcal{S}_{t}^{(1,1/2)}$ is given locally around $z=0$ by
$\mathcal{S}_{t}^{(1,1/2)}(z)$ $=\frac{z^{2}+2z-\kappa_{t}^{-1}(z)}{z(1+z)}=\mathcal{S}_{\infty}^{(1,1/2)}(z)-\frac{\kappa_{t}^{-1}(z)}{z(1+z)},$ (5)
where $\kappa_{t}$ is defined by
$\kappa_{t}(z)=\frac{(1+\sqrt{1+z})\xi_{2t}(\sqrt{1+z})+(\sqrt{1+z}-1)}{1-\xi_{2t}(\sqrt{1+z})}.$
For any time $t>0$, $\xi_{2t}$ is the inverse, in a vicinity of $u=1$, of the Herglotz transform of the spectral distribution of the free unitary Brownian motion at time $2t$ [MR1430721]:
$\xi_{2t}(u)=\frac{u-1}{u+1}e^{ut}.$
Proof.
If $\theta=1/2,\lambda=1,$ then the PDE (4) reduces to:
$\partial_{t}\mathcal{S}_{t}^{(1,1/2)}(z)$ $=(1+z)\mathcal{S}_{t}^{(1,1/2)}(z)-\frac{1+2z}{2}\left[\mathcal{S}_{t}^{(1,1/2)}(z)\right]^{2}-\frac{z(1+z)}{4}\partial_{z}\left[\mathcal{S}_{t}^{(1,1/2)}\right]^{2}(z).$ (6)
For a fixed $z$ near the origin, the characteristic curve $s\mapsto z(s)$ of the PDE (6) starting at $z$ satisfies locally (in time) the ODE
$z^{\prime}(s)=\frac{1}{2}z(s)(1+z(s))f(s),\quad z(0)=z.$ (7)
where we set:
$f(s):=\mathcal{S}_{t}^{(1,1/2)}(z(s)).$
It follows that
$f^{\prime}(s)=(1+z(s))f(t)-\frac{1+2z(s)}{2}f(s)^{2},\quad f(0)=\mathcal{S}_{0}^{(1,1/2)}(z).$
Substituting the ODE (7) into the second yields
$2\frac{z^{\prime\prime}(s)z(s)(1+z(s))-z^{\prime}(s)^{2}(1+2z(s))}{z(s)^{2}(1+z(s))^{2}}=\frac{2z^{\prime}(s)}{z(s)}-2\frac{z^{\prime}(s)^{2}(1+2z(s))}{z(s)^{2}(1+z(s))^{2}},$
which reduces to
$z^{\prime\prime}(s)=z^{\prime}(s)(1+z(s)).$
A first integration gives
$z^{\prime}(s)-z^{\prime}(0)=z^{\prime}(s)-\frac{z(1+z)}{2}f(0)=z(s)+\frac{z^{2}(s)}{2}-z-\frac{z^{2}}{2},$
or equivalently
$z^{\prime}(s)=z(s)+\frac{z^{2}(s)}{2}-z\left(1-\frac{S_{0}^{(1,1/2)}(z)}{2}\right)-\frac{z^{2}}{2}(1-S_{0}^{(1,1/2)}(z)).$
Specializing the last ODE to $\mathcal{S}_{0}^{(1,1/2)}(z)=1$, a second integration then gives (locally in time)
$\frac{z(s)+1-\sqrt{1+z}}{z(s)+1+\sqrt{1+z}}=\frac{\sqrt{z+1}-1}{\sqrt{z+1}+1}e^{\sqrt{1+z}s}=\xi_{2s}(\sqrt{1+z}),$
while equation (7) yields again:
$\mathcal{S}_{t}^{(1,1/2)}(z(s))=f(s)=\frac{z^{2}(s)+2z(s)-z}{z(s)(1+z(s))}.$
Consequently, it holds locally in time that:
$z(s)=\frac{(1+\sqrt{1+z})\xi_{2s}(\sqrt{1+z})+(\sqrt{1+z}-1)}{1-\xi_{2s}(\sqrt{1+z})}=\kappa_{s}(z).$ (8)
Now, recall from [MR1430721], Lemma 12, that for any fixed time $t$, there exists a Jordan $\Gamma_{t}$ lying in the right half-plane and containing $w=1$ where $\xi_{t}$ is a one-to-one map onto the open unit disc. As a matter of fact, there exists a neighbourhood of the origin such that for any $z$ there, the associated characteristic curve $s\mapsto z(s)$ is globally defined in time.
Finally, fix $t>0$ and assume for a while that $z\mapsto\kappa_{t}(z)$ is locally invertible around the origin. Then $z=\kappa_{t}^{-1}(z(t))$ and we finally get:
$\mathcal{S}_{t}^{(1,1/2)}(z(t))$ $=\frac{z^{2}(t)+2z(t)-\kappa_{t}^{-1}(z(t))}{z(t)(1+z(t))},$
whence the expression of $\mathcal{S}_{t}^{(1,1/2)}(z)$ follows.
Coming back to the local invertibility of $\kappa_{t}$ around the origin, observe that
$\kappa_{t}(z)=\sqrt{1+z}\frac{1+\xi_{2t}(\sqrt{1+z})}{1-\xi_{2t}(\sqrt{1+z})}-1,$
so that one only needs to check the same property for the map
$v\mapsto v\frac{1+\xi_{2t}(v)}{1-\xi_{2t}(v)},$
around $v=1$. Since the derivative of this map at $v=1$ is given by $1+2\xi_{2t}^{\prime}(1)=1+e^{t}\neq 0$, the proposition is then proved. ∎
Remark 3.3.2.
For any time $t\geq 0$, let $V_{t}$ be defined by:
$V_{t}(z)=z(1+z)\mathcal{S}_{t}^{(1,1/2)}(z),$
From the PDE (6), we readily infer
$\partial_{t}V_{t}=(1+z)V_{t}-\frac{1}{4}\partial_{z}V_{t}^{2},$
and in turn $z\mapsto\kappa_{t}^{-1}(z)=z^{2}+2z-V_{t}(z)$ satisfies
$\partial_{t}\kappa_{t}^{-1}=\frac{1}{4}\partial_{z}[\kappa_{t}^{-1}]^{2}-\frac{z^{2}+2z}{2}\partial_{z}\kappa_{t}^{-1}.$
On the other hand, the proposition yields the limit
$\lim_{t\rightarrow+\infty}\kappa_{t}^{-1}=0,$
This aligns with the fact that the Jordan domain $\Gamma_{t}$ shrinks to $w=1$ in the limit as $t\rightarrow+\infty$ and the fact that $\xi_{2t}(1)=0=\kappa_{t}(0)$ for any time $t\geq 0$.
4. Averaged characteristic polynomial of the Hermitian Jacobi Process
In this section, we proceed to the study of the averaged characteristic polynomial of the Hermitian Jacobi process. The main results are described as follows:
  • In Proposition (4.1.1), we revisit Voit’s result showing that the averaged characteristic polynomial satisfies the inverse Jacobi-heat equation for $p\wedge q>m-(1/2)$. In doing so, we prove it directly using stochastic calculus, without relying on elementary symmetric polynomials. We also emphasise that this result holds for the larger set $p\wedge q>m-1$ by virtue of Corollary 7 in [MR3296535].
  • Corollary 4.1.2 gives then an explicit expression of this polynomial in the Jacobi polynomial basis when starting at $(x-1)^{m}$, while Proposition 4.1.4 shows that all its roots remain in the interval $(0,1)$. In this respect, we would like to stress that this last result is far from being obvious from a differential equation perspective since the dynamics of the roots are highly nonlinear. Actually, it holds true in our framework since this deterministic particle system results from ’freezing’ the eigenvalues process (1).
  • Proposition 4.2.1 is the finite free analogue of Corollary 3.3. proved in [MR3071702]: when $\lambda=1,\theta=1/2$, the image of the roots of the averaged characteristic polynomial of the Hermitian Jacobi process by the transformation $x\mapsto 2\arccos(\sqrt{x})$ coincides with the roots of the Hermite unitary polynomial.
  • Proposition 4.3.1 provides an expansion of $\chi_{t}^{(r,s,m)}$ which is valid for any initial value of the eigenvalues process (1). While Proposition (4.1.1) also shows that such expansion follows from the inverse Jacobi-heat equation, the one provided by Proposition 4.3.1 relies on the so-called dual Cauchy identity satisfied by multivariate Jacobi polynomials. This identity originates in algebraic combinatorics (see [MR3075094] and references therein) and opens the way to use the mutual orthogonality of the multivariate Jacobi polynomials.
4.1. Averaged Characteristic Polynomial
Let
$F(x;u_{1},\dots,u_{m})=\prod_{i=1}^{m}(x-u_{i}),\quad x\in[0,1],$
then the averaged characteristic polynomial of $X_{t}$ is:
$\chi_{t}^{(r,s,m)}(x):=\mathbb{E}F(x;\lambda_{t}^{1},\dots,\lambda_{t}^{m}).$
In the series of papers [MR4092353], [MR4673381], [MR4781073], the roots of the characteristic polynomial of the (rescaled) $\beta$-Jacobi particle system valued in [1,1][-1,1] are studied, among others. In particular, Corollary 3.4 from [MR4092353] shows that the averaged elementary symmetric polynomials in this random system do not depend on $\beta$. Besides, Theorem 4.4 from [MR4781073] describes the dynamics of the roots of the averaged characteristic polynomial through an inverse Jacobi-heat equation. Though the $\beta$-Jacobi particle systems in $[0,1]$ and in [1,1][-1,1] are related by an affine transformation, the relation between elementary symmetric polynomials in both corresponding roots is far from being trivial. Nonetheless, one expects that the dynamics of the roots associated with the Jacobi particle system in $[0,1]$ are still governed by an inverse Jacobi-heat equation. In this regard, a direct application of Itô’s formula together with appropriate algebraic transformations shows that this is indeed true.
Proposition 4.1.1.
Let $m\geq 1$. Assume $p\wedge q>m-1$, then $(t,x)\mapsto\chi_{t}^{(r,s,m)}(x)$ solves the inverse heat equation:
$\partial_{t}\chi_{t}^{(r,s,m)}(x)=-\left\{\mathcal{L}_{x}^{(r,s)}+m(r+s+m+1)\right\}\chi_{t}^{(r,s,m)}(x),$
where
$\mathcal{L}_{x}^{(r,s)}:=x(1-x)\partial_{xx}^{2}+[(r+1)-(r+s+2)x]\partial_{x}$
is the one-dimensional Jacobi operator.
Proof.
Let
$F(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})=\prod_{i=1}^{m}(x-\lambda_{t}^{i}),\quad x\in\mathbb{R},$
be the characteristic polynomial of the Hermitian Jacobi process $X_{t}$ and recall that the eigenvalues are almost surely simple. Then, Itô’s formula entails:
$dF(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})=\textrm{Local Martingale}-F(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})\sum_{i=1}^{m}\frac{p-(p+q)\lambda_{t}^{i}}{x-\lambda_{t}^{i}}\\ -F(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})\sum_{i=1}^{m}\frac{1}{x-\lambda_{t}^{i}}\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})+\lambda_{t}^{j}(1-\lambda_{t}^{i})}{\lambda_{t}^{i}-\lambda_{t}^{j}}dt.$
Note in passing that the quadratic variation terms have zero contribution since
$\partial_{u_{i}u_{i}}F(x,u_{1},\dots,u_{m})=0,$
and since the Brownian motions $(B_{t}^{i},t\geq 0)_{1\leq i\leq m}$ are independent. Using
$\partial_{x}F(x,u_{1},\dots,u_{m})=F(x,u_{1},\dots,u_{m})\sum_{i=1}^{m}\frac{1}{x-u_{i}},$
it follows that
$F(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})\sum_{i=1}^{m}\frac{p-(p+q)\lambda_{t}^{i}}{x-\lambda_{t}^{i}}$ $=p\partial_{x}F(x,\lambda_{t})-(p+q)F(x,\lambda_{t})\sum_{i=1}^{m}\frac{\lambda^{i}_{t}}{x-\lambda_{t}^{i}}$
$=p\partial_{x}F(x,\lambda_{t})-(p+q)F(x,\lambda_{t})\sum_{i=1}^{m}(\frac{x}{x-\lambda^{i}_{t}}-1)$
$=p\partial_{x}F(x,\lambda_{t})-(p+q)(x\partial_{x}F(x,\lambda_{t})-mF(x,\lambda))$
$=\partial_{x}F(x,\lambda_{t})(p-(p+q)x)+m(p+q)F(x,\lambda_{t})$
where we used the shorthand notation $(x,\lambda_{t})$ to denote $(x,\lambda_{t}^{1},\dots,\lambda_{t}^{m})$. Furthermore, the term
$2\sum_{i=1}^{m}\frac{1}{x-\lambda_{t}^{i}}\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})+\lambda_{t}^{j}(1-\lambda_{t}^{i})}{\lambda_{t}^{i}-\lambda_{t}^{j}}$
may be symmetrised as
$\sum_{i=1}^{m}\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})+\lambda_{t}^{j}(1-\lambda_{t}^{i})}{\lambda_{t}^{i}-\lambda_{t}^{j}}\left[\frac{1}{x-\lambda_{t}^{i}}-\frac{1}{x-\lambda_{t}^{j}}\right]$ $=\sum_{i=1}^{m}\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})+\lambda_{t}^{j}(1-\lambda_{t}^{i})}{(x-\lambda_{t}^{i})(x-\lambda_{t}^{j})}$
$=2\sum_{i=1}^{m}\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})}{(x-\lambda_{t}^{i})(x-\lambda_{t}^{j})}.$
On the other hand, one readily computes:
$\partial_{xx}^{2}F(x,u_{1},\dots,u_{m})=F(x,u_{1},\dots,u_{m})\sum_{i=1}^{m}\sum_{j\neq i}\frac{1}{(x-u_{i})(x-u_{j})}$
and split, for any $1\leq i\neq j\leq m$,
$\frac{x(1-x)}{(x-u_{i})(x-u_{j})}$ $=\frac{1-x}{x-u_{j}}+\frac{u_{i}(1-x)}{(x-u_{i})(x-u_{j})}$
$=\frac{1-x}{x-u_{j}}+\frac{u_{i}(1-u_{j})}{(x-u_{i})(x-u_{j})}-\frac{u_{i}}{x-u_{i}}$
$=\frac{1-x}{x-u_{j}}+\frac{u_{i}(1-u_{j})}{(x-u_{i})(x-u_{j})}-\frac{x}{x-u_{i}}+1.$
Consequently, again with the shorthand notation $(x,u)$ for $(x,u_{1},\dots,u_{m})$, one gets:
$x(1-x)\partial_{xx}^{2}F(x,u)$ $=[m(m-1)+(m-1)(1-x)\partial_{x}-(m-1)x\partial_{x}]F(x,u)$
$+F(x,u)\sum_{i=1}^{m}\sum_{j\neq i}\frac{u_{i}(1-u_{j})}{(x-u_{i})(x-u_{j})},$
whence we deduce that
$F(x,\lambda)\sum_{i=1}^{m}$ $\sum_{j\neq i}\frac{\lambda_{t}^{i}(1-\lambda_{t}^{j})}{(x-\lambda_{t}^{i})(x-\lambda_{t}^{j})}=x(1-x)\partial_{xx}^{2}F(x,\lambda)-(m-1)[m+(1-2x)\partial_{x}]F(x,\lambda).$
Gathering all the terms, we get the following SDE for the characteristic polynomial of $X_{t}$:
$d[F(x,\lambda_{t})]=\textrm{Local Martingale}+\\ -x(1-x)\partial_{xx}^{2}F(x,\lambda_{t})-(r+1-(s+r+2)x)\partial_{x}F(x,\lambda_{t})-m(d-m+1)F(x,\lambda_{t})$ (9)
Finally, the bracket of the local martingale part is given by
$2\sum_{i=1}^{m}\lambda_{t}^{i}(1-\lambda_{t}^{i})\prod_{j\neq i}(x-\lambda_{j}),$
and is obviously bounded for any fixed $x$ in a bounded interval. Consequently, the local martingale part is a true martingale, and the inverse Jacobi-heat equation follows after taking the expectation of both sides of (9). ∎
It is known that the Jacobi operator $\mathcal{L}_{x}^{(r,s)}$ admits a complete set of eigenpolynomials given by the orthogonal Jacobi polynomials $Q_{j}^{(r,s)}(x),j\geq 0$:
$Q_{j}^{(r,s)}(x):=\frac{(r+1)_{j}}{j!}{}_{2}F_{1}(-j,r+s+j+1,r+1;x),$
where $_{2}F_{1}$ stands for the Gauss hypergeometric function. More precisely, one has:
$\mathcal{L}_{x}^{(r,s)}[Q_{j}^{(r,s)}(\cdot)](x)=-j(j+r+s+1)Q_{j}^{(r,s)}(x).$
Consequently,
$\left\{\mathcal{L}_{x}^{(r,s)}+m(d-m+1)\right\}Q_{j}^{(r,s)}(\cdot)(x)$ $=\left\{-j(j+r+s+1)+m(r+s+m+1)\right\}Q_{j}^{(r,s)}(x)$
$=(m-j)(r+s+1+m+j)Q_{j}^{(r,s)}(x).$
Proposition 4.1.1 then yields the first part of the following corollary.
Corollary 4.1.2.
With $p,q,m$ as in Proposition 4.1.1, let
$\chi_{0}^{(r,s,m)}(x)=\sum_{j=0}^{m}c_{j}^{(r,s,m)}Q_{j}^{(r,s)}(x)$
be the expansion of the averaged characteristic polynomial at $t=0$ in the Jacobi polynomial basis. Then
$\chi_{t}^{(r,s,m)}(x)=\sum_{j=0}^{m}c_{j}^{(r,s,m)}e^{-(m-j)(r+s+1+m+j)t}Q_{j}^{(r,s)}(x).$
In particular, if $\chi_{0}^{(r,s,m)}(x)=(x-1)^{m}$, then
$\chi_{t}^{(r,s,m)}(x)=(-1)^{m}m!\sum_{j=0}^{m}e^{-(m-j)(r+s+1+m+j)t}Q_{j}^{(r,s)}(x)\frac{(r+s+1+2j)\Gamma(r+s+1+j)}{\Gamma(r+s+2+m+j)}\binom{m+s}{m-j}$.
Proof.
It suffices to find the expansion of $(x-1)^{m}$ in the Jacobi polynomial basis. But this is afforded by the following formula [MR1697414]:
$\left(\frac{1-u}{2}\right)^{m}=\sum_{j=0}^{m}\frac{(-m)_{j}(\alpha+j+1)_{m-j}(\alpha+\beta+2j+1)}{(\alpha+\beta+j+1)_{m+1}}P_{j}^{(\alpha,\beta)}(u),\quad u\in[-1,1],$
where $P_{j}^{(\alpha,\beta)}$ is the $j$-th Jacobi polynomial in [1,1][-1,1]:
$P_{j}^{(\alpha,\beta)}(1-2x)=Q_{j}^{(\alpha,\beta)}(x),\quad\alpha,\beta>-1.$
Indeed, the symmetry property $P_{j}^{(\alpha,\beta)}(u)=P_{j}^{(\beta,\alpha})(-u)$ transforms the last formula into
$\left(\frac{1+u}{2}\right)^{m}=\sum_{j=0}^{m}(-1)^{j}\frac{(-m)_{j}(\alpha+j+1)_{m-j}(\alpha+\beta+2j+1)}{(\alpha+\beta+j+1)_{m+1}}P_{j}^{(\beta,\alpha)}(u).$
Substituting $u=1-2x,\beta=r,\alpha=s$ there, we obtain the expansion:
$(1-x)^{m}=m!\sum_{j=0}^{m}\frac{\Gamma(s+m+1)\Gamma(r+s+j+1)(r+s+2j+1)}{(m-j)!\Gamma(j+s+1)\Gamma(r+s+m+j+2)}Q_{j}^{(r,s)}(x)$ (10)
where we used the representation of the Pochhammer symbol valid for positive real numbers:
$(y)_{j}=\frac{\Gamma(y+j)}{\Gamma(y)},\quad y>0.$
Applying the shifted Jacobi-heat semi-group $e^{-t(\mathcal{L}_{x}^{(r,s)}+m(d-m+1))}$ to (10), we are done. ∎
Remark 4.1.3.
As $t\rightarrow+\infty$, only the term $j=m$ in the above expansion of $\chi_{t}^{(r,s,m)}(x)$ gives a non zero contribution. Accordingly, the limiting averaged characteristic polynomial
$\chi_{\infty}^{(r,s,m)}(x):=\lim_{t\rightarrow+\infty}\chi_{t}^{(r,s,m)}(x)$
is proportional to the monic Jacobi polynomial $Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$, where $k_{m}^{(r,s)}$ is the leading term of $Q_{m}^{(r,s)}(x)$. This is in agreement with the fact that the averaged characteristic polynomial of the Jacobi unitary ensemble, the weak limit of the Hermitian Jacobi process as $t\rightarrow+\infty$, is given by a Jacobi polynomial. Indeed, the eigenvalues of the JUE (Jacobi Unitary Ensemble) are given by a Selberg weight, as such its averaged characteristic polynomial is an instance of the celebrated Aomoto integral (see e.g. [MR1985318]).
For any fixed time $t\geq 0$, let $(x_{k}^{(r,s,m)}(t))_{1\leq k\leq m}$ be the root sequence of the polynomials $\chi^{(r,s,m)}_{t}$:
$\chi_{t}^{(r,s,m)}(x)=\prod_{m\geq k\geq 1}(x-x_{k}^{(r,s,m)}(t))=\mathbb{E}\left[\prod_{m\geq i\geq 1}(x-\lambda_{i}(t))\right].$ (11)
Then this sequence admits the following properties:
Proposition 4.1.4.
The roots $(x_{j}^{(r,s,m)}(t))_{1\leq j\leq m}$ of $\chi_{t}^{(r,s,m)}$ are all real. In addition, up to re-indexing, they satisfy the following ODE:
$\frac{dx_{j}^{(r,s,m)}}{dt}(t)$ $=(p-(p+q)x_{j}^{(r,s,m)}(t))$ (12)
$\hskip 56.9055pt+\sum_{k\neq j}\frac{x_{j}^{(r,s,m)}(t)(1-x_{k}^{(r,s,m)}(t))+x_{k}^{(r,s,m)}(t)(1-x_{j}^{(r,s,m)}(t))}{x_{j}^{(r,s,m)}(t)-x_{k}^{(r,s,m)}(t)},$
$=(r+1)-(r+s+2)x_{j}^{(r,s,m)}(t)$
$\hskip 56.9055pt+2x_{j}^{(r,s,m)}(t)(1-x_{j}^{(r,s,m)}(t))\sum_{k\neq j}\frac{1}{x_{j}^{(r,s,m)}(t)-x_{k}^{(r,s,m)}(t)}.$
Besides, for any time $t>0$ and any $1\leq j\leq m$, $x_{j}^{(r,s,m)}(t)\in(0,1)$.
Proof.
The first part of this proposition is due to Voit, as stated in Theorem 4.4 of [MR4781073]. Now, let’s turn to the second statement. We have
$\prod_{k=1}^{m}x_{k}^{(r,s,m)}(t)=\mathbb{E}\left[\prod_{k=1}^{m}\lambda_{k}(t)\right]>0$
since $\lambda_{k}(t)>0$ almost surely for any $1\leq k\leq m$ [MR2719370, MR3296535]. It follows that $x_{k}^{(r,s,m)}(t)\neq 0$ for any $1\leq k\leq m$. Similarly,
$e_{k}^{(r,s,m)}(x_{1}^{(r,s,m)}(t),\dots,x_{m}^{(r,s,m)}(t))=\mathbb{E}[e_{k}^{(r,s,m)}(\lambda_{1}(t),\dots,\lambda_{m}(t))]>0$
which implies that $x_{k}^{(r,s,m)}(t)>0$ for any $1\leq k\leq m$. In fact, if all the elementary symmetric polynomials on real numbers $x_{1},\dots,x_{m}$ are positive, then so do $(x_{1},\dots,x_{m})$. This is proved by setting
$J(x):=\prod_{k=1}^{m}(x+x_{k})$
and argue that if $x_{i}<0$ for some $i\in[m]$, this yields the following contradiction:
$0=J(-x_{i})=\sum_{k=1}^{m}(-x_{i})^{m-k}e_{k}(x_{1},\dots,x_{m})>0.$
To prove now that $x_{k}^{(r,s,m)}(t)<1$ for any $1\leq k\leq m$, observe that
$\prod_{k=1}^{m}(x-(1-x_{k}^{(r,s,m)}(t))=\prod_{k=1}^{m}(x_{k}^{(r,s,m)}(t)-(1-x))=\mathbb{E}\left[\prod_{k=1}^{m}(x-(1-\lambda_{k}(t))\right].$
But $(1-\lambda_{k}(t),t\geq 0)_{1\leq k\leq m}$ is a Jacobi particle system with parameters $(s,r,m)$. As a matter of fact, $1-x_{k}^{(r,s,m)}(t)>0$. ∎
4.2. Relation to the Hermite unitary polynomial
In [MR3071702], it was shown that the spectral distribution of the free Jacobi process with parameters $(1,1/2)$ in the compressed probability space coincides with the spectral distribution of
$\frac{1}{4}(Y_{2t}+Y^{\star}_{2t}+2)=\frac{1}{2}(1+\Re(Y_{2t})),$
where $(Y_{t})_{t\geq 0}$ is the free unitary Brownian motion [MR1430721]. Note in this respect that this special set of parameters corresponds at the matrix level to Hermitian Jacobi processes, which are radial parts of ‘asymptotically square’ (as $m\rightarrow+\infty$) principal minors of a unitary Brownian motion whose size is asymptotically twice the sizes of these truncations.
On the other hand, the finite analogue of the free unitary convolution, with respect to which $(Y_{t})_{t\geq 0}$ is a free Lévy process, was introduced and studied in [MR4285332]. This finite convolution is encoded by the zeroes of the Hermite unitary polynomial defined for any time $t\geq 0$ by [MR4912666] [^2] This definition differs from the original one given in [MR4285332] by the time change $t\mapsto-t(n-1)$.:
$H_{d}(z,t)=\sum_{k=0}^{d}x^{d-k}(-1)^{k}\binom{d}{k}\exp\left(-t\frac{k(d-k)}{2}\right).$
In particular, Lemma 2.1. in [MR4912666] asserts that the zeroes of $H_{d}(\cdot,t)$ lie on the unit circle and we infer from Corollary 3.35 proved in [MR4285332] that the dynamics of the corresponding angles satisfy the following ODE:
$\partial_{t}\theta_{j}(t)=\frac{1}{2}\sum_{k\neq j}\cot\left(\frac{\theta_{j}(t)-\theta_{k}(t)}{2}\right),\quad 1\leq j\leq d.$ (13)
In particular, Theorem 6.1. in [MR4781073] shows that these angles (and in turn the roots of $H_{d}(\cdot,t)$) remain distinct for all times $t>0$ even if they collapse at $t=0$. Besides, since $H_{d}(\cdot,t)$ has real coefficients, then its roots are pairwise-conjugate. Note also that z=1z=-1 is not a root of $H_{d}(\cdot,t)$ since $H_{d}(-1,t)$ is a sum of positive terms, while $z=1$ is so only when the degree $d$ is odd. Based on this discussion, it is tempting to wonder whether both root dynamics (13) and (12) are related through the transformation [^3] For sake of simplicity, we omit the dependence on the parameters.:
$x_{j}(t)=\frac{1+\cos(\theta_{j}(2t))}{2}=\cos^{2}\left(\frac{\theta_{j}(2t)}{2}\right),\quad 1\leq j\leq m,$
when $d=2m$ is even, where here $0<\theta_{1}(t)<\theta_{2}<\cdots<\theta_{m}<\pi$ are the ordered angles of the zeroes of $H_{d}$ lying in the upper-half of the unit circle. The following proposition shows that this is indeed the case:
Proposition 4.2.1.
Assume $\chi^{(-1/2,-1/2,m)}_{0}=(x-1)^{m}=H_{2m}(x,0)$. Then, for any $t\geq 0$ and any $m\geq 0$:
$H_{2m}(z,2t)=4^{m}z^{m}\chi^{(-1/2,-1/2,m)}_{t}\left(\frac{z+z^{-1}+2}{4}\right).$ (14)
Proof.
For $m=0$, the result is immediate. We turn our attention to the case $m\geq 1$. Set $\eta_{j}(t)=\cos^{2}(\theta_{j}(t)/2)$, for $1\leq j\leq m$. Then:
$\partial_{t}\eta_{j}(t)$ $=-\left[\partial_{t}\theta_{j}(t)\right]\sin(\theta_{j}(t)/2)\cos(\theta_{j}(t)/2)$
$=-\frac{1}{2}\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq 2m\end{subarray}}\cot\left(\frac{\theta_{j}(t)-\theta_{k}(t)}{2}\right)\sin(\theta_{j}(t)/2)\cos(\theta_{j}(t)/2)$
$=-\frac{1}{2}\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq 2m\end{subarray}}\frac{\cos^{2}(\theta_{j}(t)/2)\cos(\theta_{k}(t)/2)\sin(\theta_{j}(t)/2)+\sin^{2}(\theta_{j}(t)/2)\sin(\theta_{k}(t)/2)\cos(\theta_{j}(t)/2)}{\sin((\theta_{j}(t)-\theta_{k}(t))/2)}.$
Using the identity $\sin^{2}(\theta_{j}(t)/2)=1-\cos^{2}(\theta_{j}(t)/2)$, we further get:
$\partial_{t}\eta_{j}(t)$ $=-\frac{2m-1}{2}\eta_{j}(t)-\frac{1}{2}\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq 2m\end{subarray}}\frac{\sin(\theta_{k}(t)/2)\cos(\theta_{j}(t)/2)}{\sin((\theta_{j}(t)-\theta_{k}(t))/2)}$
$=-\frac{2m-1}{2}\eta_{j}(t)-\frac{1}{2}\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq 2m\end{subarray}}\left(\frac{\tan(\theta_{j}(t)/2)}{\tan(\theta_{k}(t)/2)}-1\right)^{-1}.$
Taking out the summand corresponding to k=2mj+1k=2m-j+1 and remembering that the angles come into opposite pairs, the last sum splits into:
$\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq 2m\end{subarray}}\left(\frac{\tan(\theta_{j}(t)/2)}{\tan(\theta_{k}(t)/2)}-1\right)^{-1}$ $=-\frac{1}{2}+\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\left(\frac{\tan(\theta_{j}(t)/2)}{\tan(\theta_{k}(t)/2)}-1\right)^{-1}-\left(\frac{\tan(\theta_{j}(t)/2)}{\tan(\theta_{k}(t)/2)}+1\right)^{-1}$
$=-\frac{1}{2}+2\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\tan^{2}(\theta_{k}(t)/2)}{{\tan^{2}(\theta_{j}(t)/2)}-{\tan^{2}(\theta_{k}(t)/2)}}$
$=-\frac{1}{2}+2\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\cos^{2}(\theta_{j}(t)/2)-\cos^{2}(\theta_{j}(t)/2)\cos^{2}(\theta_{k}(t)/2)}{{\cos^{2}(\theta_{k}(t)/2)}-{\cos^{2}(\theta_{j}(t)/2)}}.$
As result, for any $1\leq j\leq m$, the map $\eta_{j}$ satisfies the ODE:
$\partial_{t}\eta_{j}(t)$ $=-\frac{2m-1}{2}\eta_{j}(t)+\frac{1}{4}-\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\cos^{2}(\theta_{j}(t)/2)-\cos^{2}(\theta_{j}(t)/2)\cos^{2}(\theta_{k}(t)/2)}{\cos^{2}(\theta_{k}(t)/2)-\cos^{2}(\theta_{j}(t)/2)}$
$=-\frac{2m-1}{2}\eta_{j}(t)+\frac{1}{4}+\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\eta_{j}(t)(1-\eta_{k}(t))}{\eta_{j}(t)-\eta_{k}(t)}$
Equivalently, $\tilde{\eta}_{j}(t):=\eta_{j}(2t)$ satisfies:
$\partial_{t}\tilde{\eta_{j}}(t)$ $=-(2m-1)\tilde{\eta}_{j}(t)+\frac{1}{2}+2\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\tilde{\eta_{j}}(t)(1-\tilde{\eta}_{k}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)}$
Finally, notice that
$2\frac{\tilde{\eta}_{j}(t)(1-\tilde{\eta}_{k}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)}$ $=\frac{\tilde{\eta}_{j}(t)(1-\tilde{\eta}_{k}(t))+\tilde{\eta}_{k}(t)(1-\tilde{\eta}_{j}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)}+\frac{\tilde{\eta}_{j}(t)(1-\tilde{\eta}_{k}(t))-\tilde{\eta}_{k}(t)(1-\tilde{\eta}_{j}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)}$
$=1+\frac{\tilde{\eta}_{j}(t)(1-\tilde{\eta}_{k}(t))+\tilde{\eta}_{k}(t)(1-\tilde{\eta}_{j}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)},$
we end up with:
$\partial_{t}\tilde{\eta_{j}}(t)$ $=\left(m-\frac{1}{2}\right)-(2m-1)\tilde{\eta}_{j}(t)+\sum_{\begin{subarray}{c}k\neq j\\ 1\leq k\leq m\end{subarray}}\frac{\tilde{\eta}_{j}(t)(1-\tilde{\eta}_{k}(t))+\tilde{\eta}_{k}(t)(1-\tilde{\eta}_{j}(t))}{\tilde{\eta}_{j}(t)-\tilde{\eta}_{k}(t)}.$
This is the ODE (12) with p=q=m(1/2)p=q=m-(1/2). But Theorem 1.1. in [MR4751304] shows (after performing an affine variable change) that (12) admits a unique solution for any p,q>m1p,q>m-1 which remains in the open domain $\{0<x_{m}<\cdots<x_{1}<1\}$ at any time $t>0$. It follows that
$\cos^{2}\left(\frac{\theta_{j}(2t)}{2}\right)=\tilde{\eta_{j}}(t)=x_{j}(t)$
for any $1\leq j\leq m$ and any $t\geq 0$, therefore
$H_{2m}(z,2t)$ $=\prod_{j=1}^{m}|z-e^{i\theta_{j}(2t)}|^{2}$
$=\prod_{j=1}^{m}(z^{2}-2z\cos(\theta_{j}(2t))+1)$
$=\prod_{j=1}^{m}(z^{2}-2z(2x_{j}(t)-1)+1)=(4z)^{m}\chi^{(-1/2,-1/2,m)}_{t}\left(\frac{z+z^{-1}+2}{4}\right).$
Remark 4.2.2.
We can infer a relation between the Hermitian and Jacobi heat generators. Let $\mathbb{R}_{[0,1]}[x]_{m}$ be the set of monic polynomials with degree $m$ and roots in $[0,1]$ and define:
$\varphi_{m}\colon\mathbb{R}_{[0,1]}[x]_{m}\to\mathbb{R}[z]_{2m}\quad P\mapsto 4^{m}z^{m}P(\frac{1}{4}(z+z^{-1}+2)).$
Recall the definition of the Hermitian heat generators $\mathcal{L}^{(2m)}_{A}$ and the relation with the Hermitian Unitary polynomial:
$H_{2m}(z,2t)=\exp(-t(z\partial_{z}){(2m-z\partial_{z})})\{(z-1)^{2m}\}=\exp(\mathcal{L}_{A}^{2m})\{(z-1)^{2m}\}.$
Then,
$\varphi\circ\mathcal{L}^{(-1/2,-1/2,m)}=\mathcal{L}_{A}^{(2m)}\circ\varphi.$ (15)
The polynomial $\varphi_{m}(P)$ has real coefficients, hence its roots are two-by-two conjugate. Moreover, $z^{2m}\varphi(\frac{1}{z})=\varphi(z)$ (it is invariant by inversion). Hence, the set of roots of $\varphi_{m}(P)$ is invariant under inversion and conjugation. Moreover, if $z_{0}$ of $\varphi_{m}(P)$ then $\Im(z_{0}+z_{0}^{-1})=0$ since $\frac{1}{4}(z_{0}+z_{0}^{-1}+2)$ is root of $P$. Hence, $|z_{0}|-|z_{0}|^{-1}=0$ and $|z_{0}|=1$. Thus $\varphi_{m}(P)$ has roots on the unit circle and they are pair-wise conjugate. The dynamic of the angles of the roots of $\exp(t\mathcal{L}_{A}^{(2m)})\{\varphi_{m}(P)\}$ is prescribed by (13). Equality (15) follows from the same reasoning as exposed in the proof of Proposition 4.2.1.
4.3. A more general expansion
We derive a more general expansion of the averaged characteristic polynomial $\chi_{t}^{(r,s,m)}$ which is valid for any initial value $w=\lambda_{0}$ of the eigenvalues process. Our main ingredients are the heat kernel of the latter and the dual Cauchy identity satisfied by the (symmetric) multivariate Jacobi polynomials recalled below. These polynomials are mutually orthogonal with respect to the unitary Selberg weight. This property is not satisfied by any orthogonal set of multivariable polynomials since it requires the orthogonality of any two polynomials corresponding to different partitions. For the sake of completeness, we provide a brief reminder of the key facts and results necessary to prove the expansion below. We refer the reader to the paper [MR2719370] for more details.
Let
$\tau=(\tau_{1}\geq\tau_{2}\geq...\geq\tau_{m}\geq 0)$
be a partition of length at most $m$ and let $(\tilde{Q}_{j}^{(r,s)})_{j\geq 0}$ be the sequence of orthonormal Jacobi polynomials with respect to the beta weight:
$u^{r}(1-u)^{s}{\bf 1}_{[0,1]}(u).$
These are given by
[…]
Then the orthonormal multivariate Jacobi polynomial corresponding to $\tau$ is defined by:
$\tilde{Q}_{\tau}^{(r,s,m)}(y_{1},...,y_{m}):=\frac{\det(\tilde{Q}_{\tau_{i}-i+m}^{(r,s)}(y_{j}))_{1\leq i,j\leq m}}{V(y_{1},...,y_{m})},$
where
$V(y_{1},...,y_{m}):=\prod_{1\leq i<j\leq m}(y_{i}-y_{j}),$
is the Vandermonde determinant. If the coordinates $(w_{1},...,w_{m})$ overlap, this definition still makes sense by either applying L’Hôpital’s rule or equivalently by using the expansion of $\tilde{Q}_{\tau}^{(r,s,m)}$ in the Schur polynomial basis.
The multivariable Jacobi polynomials are symmetric (invariant under permutations) and satisfy the remarkable property of being mutually orthonormal with respect to the unitary Selberg weight:
$W^{(r,s,m)}(y_{1},\dots,y_{m}):=[V(y_{1},\dots,y_{m})]^{2}\prod_{i=1}^{m}y_{i}^{r}(1-y_{i})^{s}{\bf 1}_{0<y_{m}<...<y_{1}<1},\quad r,s>-1.$
Actually, for any two different partitions $\tau$ and $\kappa$, one has:
$\int Q_{\tau}^{(r,s,m)}(y_{1},...,y_{m})Q_{\kappa}^{(r,s,m)}(y_{1},...,y_{m})W^{(r,s,m)}(y_{1},\dots,y_{m})dy_{1}\cdots dy_{m}=0,$
as one readily checks using Andreief’s identity.
Replacing $\tilde{Q}_{j}^{(r,s)}$ with $Q_{j}^{(r,s)}$, we get the orthogonal multivariate Jacobi polynomials $(Q_{\tau}^{(r,s,m)})_{\tau}$ in $0,1]^{m}$, and performing further the variable change $y\mapsto 1-2y$, one gets the orthogonal multivariate Jacobi polynomials $(P_{\tau}^{(r,s,m)})_{\tau}$ in $1,1]^{m}$.
Now, recall that the semi-group density of the eigenvalues process of $X_{t}$ starting at $w$ admits the following absolutely-convergent expansion:
$G_{t}^{r,s,m}(w,y):=\sum_{\tau}e^{-\nu_{\tau}t}\tilde{Q}_{\tau}^{(r,s,m)}(w)\tilde{Q}_{\tau}^{r,s,m}(y)W^{r,s,m}(y_{1},\dots,y_{m}),\quad r,s>-1,$
where
$\nu_{\tau}:=\sum_{i=1}^{m}\tau_{i}(\tau_{i}+r+s+1+2(m-i)).$
Now, we are ready to prove the following proposition:
Proposition 4.3.1.
For any $w\in[0,1]^{m}$,
$\chi_{t}^{(r,s,m)}(x)=\frac{1}{(-2)^{m(m+1)/2}}\sum_{j=0}^{m}(-1)^{m-j}e^{-\nu_{(1^{m-j})}t}\frac{Q_{(1^{m-j})}^{(r,s,m)}(w)}{k_{1^{m-j}}^{(r,s,m)}}\frac{Q_{j}^{(r,s)}(x)}{k_{j}^{(r,s)}},$
where $1^{m-j}$ is the partition with only (mj)(m-j) ones, $k_{1^{m-j}}^{(r,s,m)}$ is the leading coefficient of $P_{(1^{m-j})}^{(r,s,m)}$ and $k_{j}^{(r,s)}=k_{j}^{(r,s,1)}$ is the leading coefficient of $P_{j}^{(r,s)}:=P_{j}^{(r,s,1)}$.
Proof.
Recall from [MR3075094] the dual Cauchy-identity:
$\prod_{i=1}^{N}\prod_{j=1}^{K}(u_{i}+v_{j})=\sum_{\begin{subarray}{c}\lambda=(\lambda_{1}\geq\dots\lambda_{N}\geq 0)\\ \lambda_{1}\leq K\end{subarray}}\frac{P_{\mu}^{(s,r,m)}(v_{1},\dots,v_{K})P_{\lambda}^{(r,s,m)}(u_{1},\dots,u_{N})}{k_{\mu}^{(s,r,m)}k_{\lambda}^{(r,s,m)}},\quad u_{i},v_{j}\in[-1,1],$
where
$\mu=(N-\lambda^{\prime}_{K},\dots,N-\lambda^{\prime}_{1}),$
$\lambda^{\prime}=(\lambda^{\prime}_{1}\geq\dots\geq\lambda^{\prime}_{K})$ is the conjugate partition of $\lambda$ and $k_{\lambda}^{(r,s,m)}$ is the leading coefficients $P_{\lambda}^{(r,s,m)}$. Note in passing that the representation
$P_{j}^{(r,s)}(u)=\frac{(r+1)_{j}}{j!}{}_{2}F_{1}\left(-j,r+s+j+1,r+1;\frac{1-u}{2}\right)$
shows that
$k_{j}^{(r,s,1)}=\frac{(r+s+1+j)_{j}}{j!2^{j}},$
and in turn, the determinantal form of $P_{j}^{(r,s,m)}$ entails
$k_{\lambda}^{(r,s,m)}=\prod_{j=1}^{m}k_{\lambda_{j}-j+m}^{(r,s,1)},$
for any partition $\lambda$.
We specialise this formula by putting
$N=1,\quad K=m,\quad u_{1}=1-2x,\quad v_{j}=2y_{j}-1.$
Then $\lambda=\lambda_{1}\in\{0,\dots,m\}$ is a row so that $\lambda^{\prime}=1^{\lambda_{1}}$ is a column and in turn $\mu=1^{m-\lambda_{1}}$ is a column as well. Using the mirror property satisfied by the Jacobi polynomials in [1,1][-1,1], we get
$(-2)^{m}F(x,y_{1},\dots,y_{m})$ $=\sum_{j=0}^{m}\frac{P_{1^{m-j}}^{(s,r,m)}(2y_{1}-1,\dots,2y_{m}-1)P_{j}^{(r,s,1)}(1-2x)}{k_{1^{m-j}}^{(r,s,m)}k_{j}^{(r,s,1)}}$
$=\sum_{j=0}^{m}(-1)^{m-j}\frac{P_{1^{m-j}}^{(r,s,m)}(1-2y_{1},\dots,1-2y_{m})P_{j}^{(r,s)}(1-2x)}{k_{1^{m-j}}^{(r,s,m)}k_{j}^{(r,s,1)}}$
$=\sum_{j=0}^{m}(-1)^{m-j}\frac{Q_{1^{m-j}}^{(r,s,m)}(y_{1},\dots,y_{m})Q_{j}^{(r,s)}(x)}{(-2)^{m(m-1)/2}k_{1^{m-j}}^{(r,s,m)}k_{j}^{(r,s,1)}}.$
Now, we can write the heat kernel $G_{t}^{(r,s,m)}$ as:
[…]
then appeal to the mutual orthogonality of $(Q_{\tau}^{(r,s,m)})_{\tau}$ to compute the integral:
$\chi_{t}^{(r,s,m)}(x)=\int F(x,y_{1},\dots,y_{m})G_{t}^{(r,s,m)}(w,y)dy_{1}\cdots dy_{m}.$
Doing so only leaves the partitions $\tau=1^{m-j},0\leq j\leq m$ whence the sought expansion follows. ∎
Remark 4.3.2.
Proposition 7.1 in [MR3075094] gives an explicit expression of $Q_{\tau}^{(r,s,m)}(1^{m})$ as a ratio of products of Gamma functions. After lengthy (but easy) computations, one retrieves the second statement of Corollary 4.1.2. Moreover, if $(w_{i}=z_{i}^{(r,s)})_{1\leq i\leq m}$ are the zeroes of the Jacobi polynomials $Q_{m}^{(r,s)}$, then $Q_{(1^{m-j})}^{(r,s,m)}(w)=0$ for all $0\leq j\leq m-1$ since $Q_{\tau_{1+m-1}}^{(r,s)}(z_{i})=0$ for any $1\leq i\leq m$. Consequently, $\chi_{t}^{(r,s,m)}(x)=Q_{m}^{(r,s)}(x)/k_{m}^{(r,s)}$ for any $t\geq 0$ and agrees with the fact that $(z_{i})_{1\leq i\leq m}$ is the stationary solution of (12) (see Proposition 7 in [MR4673381]).
5. Frozen Hermitian Jacobi process and finite free probability
In this section, we study the Frozen Jacobi process and its finite S transform. The main results are
  • The convergence of the counting measure of the roots of $\chi_{t}^{(r,s,m)}$ to the distribution of the free Jacobi process at time $t\geq 0$. Stated in Corollary 5.1.2, it follows from the differential system derived in Proposition 5.1.1 and satisfied by the moment sequence of this counting measure, together with their relative compactness.
  • Theorem 5.2.1 where we establish a general result, of independent interest, regarding the convergence of the finite differences of the finite free $S$ transform.
  • Corollary 5.3.2 and Theorem 5.3.3 where the PDE in Theorem 3.1.1 is derived as a limit of an “ODE with finite differences” for the finite free $T$ transform.
5.1. High-dimensional regime
We start by proving that in the high-dimensional regime afforded by (2), the counting measure
$\mu_{t}^{(r,s,m)}=\frac{1}{m}\sum_{i=1}^{m}\delta_{x_{i}(t)}$
of the roots of $\chi_{t}^{(r,s,m)}$ converges to the spectral distribution of the free Jacobi process at time $t>0$, provided that the convergence holds at $t=0$. For any time $t\geq 0$, let
$m_{\ell}^{(r,s,m)}(t)(t)=\mu^{(r,s,m)}_{t}(x^{\ell}),\quad\ell\geq 0,$
be the moment sequence of $\mu_{t}^{(r,s,m)}$.
Proposition 5.1.1.
For any $r,s$ and $m\geq 1$ an integer, the trajectory of the infinite sequence of moments $(m_{\ell}^{(r,s,m)}(t)(t))_{\ell\geq 0}$ is the solution of the following lower triangular differential system :
$\frac{d}{dt}m_{1}^{(r,s,m)}(t)=p-(p+q)m_{1}^{(r,s,m)}(t),$
and for any $\ell\geq 2$ :
$\frac{d}{dt}m_{\ell}^{(r,s,m)}(t)(t)$ $=-\ell(p+q-\ell+1)m_{\ell}^{(r,s,m)}(t)(t)$ (16)
$\hskip 28.45274pt+\ell(p-\ell+1)m_{\ell-1}^{(r,s,m)}(t)+m\ell\sum_{k=0}^{\ell-2}(m_{k}^{(r,s,m)}(t)(t)-m_{k+1}^{(r,s,m)}(t)(t))m_{\ell-1-k}^{(r,s,m)}(t)(t).$
Proof.
Appealing to the ODE (12), we derive
$\frac{d}{dt}m_{\ell}^{(r,s,m)}(t)(t)$ $=\frac{\ell}{m}\sum_{i=1}^{m}(x^{i}(t))^{\ell-1}\frac{d}{dt}x^{i}(t)$
$=\frac{\ell}{m}\bigg(mpm_{\ell-1}^{(r,s,m)}(t)-m(p+q)m_{\ell}^{(r,s,m)}(t)(t)$
$\hskip 56.9055pt+\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell}(1-x^{j}(t))+(x^{i}(t))^{\ell-1}x^{j}(t)(1-x^{i}(t))}{x^{i}(t)-x^{j}(t)}\bigg).$
Now, we expand
$\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell}(1-x^{j}(t))+(x^{i}(t))^{\ell-1}x^{j}(t)(1-x^{i}(t))}{x^{i}(t)-x^{j}(t)}$
$\hskip 142.26378pt=\sum_{i=1}^{m}\sum_{j\neq i}\frac{-2(x^{i}(t))^{\ell}x^{j}(t)}{x^{i}(t)-x^{j}(t)}+\frac{(x^{i}(t))^{\ell}+(x^{i}(t))^{\ell-1}x^{j}(t)}{x^{i}(t)-x^{j}(t)},$
and make the first double sum for $l\geq 2$ symmetric as:
$-2\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell-1}(x^{i}(t)x^{j}(t))}{x^{i}_{t}-x^{j}(t)}$ $=-2\sum_{i=1}^{m}\sum_{j\neq i}\frac{x^{i}(t)x^{j}(t)[(x^{i}(t))^{\ell-1}-(x^{j}(t))^{\ell-1}]}{x^{i}_{t}-x^{j}(t)}$
$=-\sum_{k=0}^{\ell-2}\sum_{i=1}^{m}\sum_{j\neq i}x^{i}(t)x^{j}(t)(x^{i}(t))^{k}(x^{j}(t))^{\ell-2-k}$
$=-\sum_{k=0}^{\ell-2}\sum_{i=1}^{m}\sum_{j\neq i}(x^{i}(t))^{k+1}(x^{j}(t))^{\ell-1-k}$
$=-m^{2}\sum_{k=0}^{\ell-2}m_{k+1}^{(r,s,m)}(t)m_{\ell-1-k}^{(r,s,m)}(t)+m(\ell-1)m_{\ell}^{(r,s,m)}(t).$
Proceeding in the same way, the second term may be written as
$\sum_{i=1}^{m}\sum_{j\neq i}\frac{(x^{i}(t))^{\ell}+(x^{i}(t))^{\ell-1}x^{j}(t)}{x^{i}_{t}-x^{j}(t)}$ $=m^{2}\sum_{k=0}^{\ell-2}m_{k}^{(r,s,m)}(t)m_{\ell-1-k}^{(r,s,m)}(t)-m(\ell-1)m_{\ell-1}^{(r,s,m)}.$
Gathering all the terms, we get (16). The ODE satisfied by $m_{1}^{(r,s,m)}$ is derived exactly along the same lines noticing that the double sum is empty. ∎
We now proceed to the aforementioned convergence result. To this end, recall from [MR2384475] that the moments
$m_{\ell}(t):=\frac{\tau(X_{t}^{\ell})}{\tau(P)},\quad\ell\geq 1,\quad m_{0}(t)=1,$
of the free Jacobi process
$X_{t}=PU_{t}QU_{t}^{\star}P,$
viewed as an operator in the compressed algebra $(P\mathscr{A}P,\tau/\tau(P))$, satisfy the differential system:
$\frac{d}{dt}m_{\ell}(t)=-\ell m_{\ell}(t)+\ell\theta m_{\ell-1}(t)+\ell\lambda\theta\sum_{j=0}^{\ell-2}m_{\ell-j-1}(t)[m_{j}(t)-m_{j+1}(t)],$ (17)
where $\tau(P)=\lambda\theta\in(0,1],\tau(Q)=\theta\in(0,1]$.
Corollary 5.1.2.
Assume that for any $\ell\geq 1$, $m_{\ell}^{(r(m),s(m),m)}(0)$ converges as $m$ goes to infinity in the regime described in (2). Then, for any time $t>0$,
$\tilde{m}_{\ell}^{(r(m),s(m),m)}(t):=m_{\ell}^{(r(m),s(m),m)}(t)\left(\frac{t}{d(m)}\right)$
converges to $m_{\ell}(t)$ as well.
Proof.
Given that for all $m$, $\mu^{(r,s,m)}_{t/d(m)}$ has support in $(0,1)$, the family formed by these measures is tight. It is then sufficient to prove that there is a unique limit point with respect to the weak topology. However, again by the compactness of the support, we find it better to prove the same statement for the moments. Indeed, for any $\ell\geq 0$, the family $\{\tilde{m}_{\ell}^{(r(m),s(m),m)}(t),\,t\geq 0\}$ is uniformly bounded in $m$. Moreover, we readily infer from (16) that for any $\ell\geq 1$,
$\frac{d}{dt}\tilde{m}_{\ell}^{(r,s,m)}(t)=\ell\frac{p-\ell+1}{d}\tilde{m}_{\ell-1}^{(r,s,m)}-\ell\frac{d-\ell+1}{d}\tilde{m}_{\ell}^{(r,s,m)}+\frac{m}{d}\ell\sum_{k=0}^{\ell-2}(\tilde{m}_{k}^{(r,s,m)}-\tilde{m}_{k+1}^{(r,s,m)})\tilde{m}_{\ell-1-k}^{(r,s,m)}.$ (18)
Since $\ell$ is fixed, then the right-hand side of (18) is uniformly bounded for sufficiently large $m,p(m),d(m)$. Consequently, $\|\partial_{t}\tilde{m}_{\ell}^{(r(m),s(m),m)}\|_{\infty}$ is so whence equicontinuity follows. By Arzelà-Ascoli Theorem, the moments sequence $\{\tilde{m}^{(r(m),s(m),m)}\}_{m}$ converges locally uniformly in $t$ along a subsequence and in turn so does $d\tilde{m}^{(r(m),s(m),m)}/dt$. Passing to the limit in (18) and using induction on $\ell$, we conclude that any limiting moment sequence satisfies (17) for any $\ell$. Since the latter admits a unique solution for any fixed intial data at $t=0$, there is one and only one accumulation point for $\tilde{m}_{\ell}^{(r(m),s(m),m)}(t)$ with prescribed initial data. The proposition is proved. ∎
5.2. Finite free probability and finite differences of the finite free $S$ transforms
The goal of this section is to propose another derivation of the PDE satisfied by the free $S$ transform of the Jacobi process, see Theorem 3.1.1, by using finite free probability and the finite free $S$ transform. This derivation paves the way to study more general particles systems, beyond the compact case; such as the frozen Heckman-Opdam process of type BC [MR4406927]. It requires a technical result regarding the asymptotics of the discrete derivatives of the finite free $S$ transform, which is proven in the following paragraphs. The reader will find all relevant definitions used in this section in the series of papers [arizmendi2018cumulants, arizmendi2018cumulants, MR4408504]. For brevity, only the basic ones are recalled below.
5.2.1. The finite free $S$ and $T$ transforms
Let $p$ be a monic polynomial with degree $m$:
$p_{m}(x)=\sum_{k=0}^{m}(-1)^{k}e^{(m)}_{k}(p)x^{m-k},\quad e^{(m)}_{0}(p)=1.$
Let $r$ the multiplicity of 0 in $p_{m}$ and assume that $p_{m}$ has only positive real roots. Then the finite free $S$ transform of $p_{m}$, denoted hereafter by $S^{(m)}_{p_{m}}$, is the function defined on the set of points $\{-k/m,k\in\{1,m-r\}\}$ defined by [arizmendi2024s]:
${S}_{p_{m}}^{(m)}\left(-\frac{k}{m}\right):=\frac{m-k+1}{k}\frac{e^{(m)}_{k-1}(p)}{e^{(m)}_{k}(p)},\quad k\in\{1,m-r\}.$
It will also be convenient to introduce the finite free $T$ transform of $p_{m}$ : it is the piecewise right-continuous function defined on the open interval $(0,1)$ by
$T_{p_{m}}^{(m)}(v)=\begin{cases}0,&v\in(0,\frac{r}{m})\\ \displaystyle\frac{m-k+1}{k}\frac{e^{(m)}_{m-k+1}(p)}{e^{(m)}_{m-k}(p)},&v\in[\frac{k-1}{m},\frac{k}{m}),\quad k=r+1,\ldots,m.\end{cases}$
Note in passing that while the finite free S transform is defined on a finite set of real numbers, the finite free T transform is defined in the open interval $(0,1)$. This difference stems from the fact that the definition of the latter is motivated by the finite free analogue of the central limit Theorem for the free multiplicative convolution (see Remark 6.4 in [arizmendi2024s]). Nonetheless, for any $k\in\{1,\dots,m-r\},$
$T_{p_{m}}^{(m)}\left(\frac{m-k}{m}\right)=\frac{1}{\displaystyle{S}_{p_{m}}^{(m)}\left(-\frac{k}{m}\right)}.$ (19)
5.2.2. Convergence of the finite differences
We let $\nabla^{(m)}$ be the operator of finite right-differentiation with step $1/m$ acting on functions of $\mathbb{R}$:
$\nabla^{(m)}\colon\mathbb{R}^{\mathbb{R}}\to\mathbb{R}^{\mathbb{R}},\quad\nabla^{(m)}g(v):=m\left(g\left(v+\frac{1}{m}\right)-g(v)\right),\quad v\in\mathbb{R}.$
Likewise, the discrete left-derivative of $g$ is
$\tilde{\nabla}^{(m)}\colon\mathbb{R}^{\mathbb{R}}\to\mathbb{R}^{\mathbb{R}},\quad\tilde{\nabla}^{(m)}g(v)=m\left(g(v)-g\left(v-\frac{1}{m}\right)\right),\quad v\in\mathbb{R}$
Most of the following statements will be formulated by using the right-derivative operator $\nabla^{(m)}$, although all have their counterparts valid for the left derivative operator $\tilde{\nabla}$.
Before stating the next Theorem, recall that if $(p_{m})_{m\geq 1}$ is a sequence of polynomials with increasing degree $m$, we denote by $\mu\llbracket p_{m}\rrbracket$ the counting measure of its roots :
$\mu\llbracket p_{m}\rrbracket:=\frac{1}{m}\sum_{z\,:\,P(z)=0}\delta_{z}.$
If $\mu\llbracket p_{m}\rrbracket$ converges weakly to some probability measure $\mu$, we say that $(p_{m})_{m\geq 1}$ is a converging sequence. From [arizmendi2024s] (see Theorem 1.1. there), if for any $m\geq 1$, $p_{m}$ has only positive roots and if $\mu\neq\delta_{0}$, then $\mu\llbracket p_{m}\rrbracket$ converges to $\mu$ if and only if $(S^{(m)}_{p_{m}})_{m}$ converges to the free $S$-transform $S_{\mu}$ of $\mu$ in the following sense: for any $v\in(0,1-\mu\{0\})$ and any sequence $(k_{m})_{m\geq 1}$ such that
$1\leq k_{m}\leq m,\quad\lim_{m\rightarrow+\infty}\frac{k_{m}}{m}=t,$
one has:
$\lim_{m\rightarrow+\infty}{S}_{p_{m}}^{(m)}\left(-\frac{k_{m}}{m}\right)=S_{\mu}(-t).$
We will also make use of the following interpolation over the interval $(0,1)$ of the finite free $S$-transform of the polynomials $p_{m}$:
$\mathcal{S}^{(m)}_{p_{m}}\colon(0,1)\to\mathbb{R},\quad\mathcal{S}^{(m)}_{p_{m}}(v)=S^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil}{m}),\quad v\in\left(0,1-\frac{r_{m}}{m}\right),$
where we recall that $\lceil{\cdot}\rceil$ is the ceiling function and $r_{m}$ is the multiplicity of 0 in $p_{m}$ (not to be confused with r(m)=p(m)m)r(m)=p(m)-m). In this regard, we again emphasize that this interpolated S transform $\mathcal{S}^{(m)}_{p_{m}}$ is only defined on the interval $(0,1-r_{m}/m)$, while $T^{(m)}_{p_{m}}$ is defined on the whole interval $(0,1)$. The relation (19) becomes
$T^{(m)}_{p_{m}}(1-v)=\frac{1}{\mathcal{S}^{(m)}_{p_{m}}(v)},\quad v\in\left(0,1-\frac{r_{m}}{m}\right).$
Theorem 5.2.1.
Let $(p_{m})_{m\geq 1}$ be a converging sequence of monic polynomials with positive roots and increasing degrees $m$. Let $\mu$ be the limiting measure :
$\mu\llbracket p_{m}\rrbracket\underset{m,+\infty}{\longrightarrow}\mu,$
and assume that $\mu\neq\delta_{0}$. Then,
$[\nabla^{(m)}](\mathcal{S}^{(m)}_{p_{m}})(v)=-\partial_{v}{S}_{\mu}(-v)+o(1),\quad m\rightarrow+\infty,$
locally uniformly on $v\in(0,1-\mu(\{0\})$.
Remark 5.2.2.
From the statement of Theorem 5.2.1, it is expected that the iterated discrete derivatives of the finite $T$-transform converge to the iterated derivatives of the $T$ transform. We will not prove this assertion, a possible strategy would be to use the same arguments as exposed in the proof of Theorem 5.2.1 in a recursive way.
Corollary 5.2.3.
Continuing the setting of Theorem 5.2.1, if $\mu$ is the limit measure of the $p_{m}$, then for any $v\in(\mu{(\{0\})},1)$,
$\nabla^{(m)}](T^{(m)}_{p_{m}})(v)=\partial_{v}T_{\mu}(v-1)+o(1)$
as $m$ goes to infinity.
Proof.
Given that for any $v\in(\mu(\{0\}),1)$:
$T_{p_{m}}^{(m)}(v)=\frac{1}{\mathcal{S}^{(m)}_{p_{m}}(1-v)},$
we infer
$\nabla^{(m)}T^{(m)}_{p_{m}}(v)$ $=-m\frac{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})-\mathcal{S}^{(m)}_{p_{m}}(1-v)}{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})\mathcal{S}^{(m)}_{p_{m}}(1-v)}$
$=\frac{\tilde{\nabla}^{(m)}\mathcal{S}^{(m)}_{p_{m}}(1-v)}{\mathcal{S}^{(m)}_{p_{m}}(1-v-\frac{1}{m})\mathcal{S}^{(m)}_{p_{m}}(1-v)}\rightarrow_{m,+\infty}-\partial_{v}{S}_{\mu}(v-1)/{S}_{\mu}(v-1)^{2}$
where we have use the local uniform convergence of $\mathcal{S}^{(m)}_{p_{m}}$ to $\mathcal{S}_{\mu}$ (as an application of the second Dini’s theorem, since $S_{\mu}$ is monotone non-increasing) and 5.2.1 for the left derivative operator $\tilde{\nabla}^{(m)}$. Since
$\partial_{v}{S}_{\mu}(v-1)/{S}_{\mu}(v-1)^{2}=\partial_{v}T_{\mu}(v-1),$
the corollary is proved. ∎
Remark 5.2.4.
Albeit the pointwise convergence of the finite free $T$ transform $T^{(m)}_{p_{m}}$ holds over the interval $(0,1)$ [arizmendi2024s, Theorem 9.1], the convergence of its discrete derivatives is stated only over the interval $(\mu(\{0\}),1)$. The behaviour of the discrete derivatives over $(0,\mu(\{0\}))$ remains unclear.
Proof of Theorem 5.2.1.
We closely follow the line of arguments presented in [arizmendi2024s] to prove Propositions 7.2, 7.3 and 7.8 based on the following representation of the free $S$ transform. Let $v\in(0,1)$, we denote by ${\rm Dil}_{v}(\mu)$ the dilation by $v$ of the probability measure $\mu$, in particular:
$\int x^{\ell}{\rm Dil}_{v}(\mu)(dx)=v^{\ell}\int x^{\ell}\mu(dx).$
Recall also the definition of fractional additive convolution power [shlyakhtenko2020fractional] in terms of the free cumulants. If $\mu$ is compactly supported, then $\mu^{\boxplus(1/v)}$ is the unique compactly supported probability measure whose free cumulants are the $(1/v)$-multiples of those of $\mu$. With these definitions [belinschi2008remarkable]:
$S_{\mu}(-v)=-G_{{\rm Dil}_{v}(\mu^{\boxplus({1}/{v})})}(0),\quad v\in(0,1).$ (20)
where $G$ denotes the Cauchy transform. Finally, recall the definition of the operators $\partial_{k|m}$ acting in monic polynomials with degree $m$:
$\partial^{k|m}p_{m}=\frac{1}{(m)_{m-k}}p^{(m-k)},\quad 0\leq k\leq m.$ (21)
Case 1: We make the additional assumption that there exists $\varepsilon>0$ and $\eta>0$ such that all the roots of the polynomials $p_{m}$ are contained in $\varepsilon,\eta)$.
Under this assumption, the convergence $\mu\llbracket p_{m}\rrbracket\to\mu$ is equivalent to the convergence of all the moments of $\mu\llbracket p_{m}\rrbracket\to\mu$ toward those of $\mu$ and further to the convergence of the finite free cumulants of $\mu\llbracket p_{m}\rrbracket\to\mu$ toward those of $\mu$. This is out of the scope of the present article to give an account about finite free cumulants, we refere the reader to [arizmendi2018cumulants].
We begin by relating the finite free $S$-transform to the Cauchy transform of the counting measure of the roots of derivatives of $p_{m}$, (see [arizmendi2024s], Lemma 6.11 2). Let $v\in(0,1)$.
$m(\mathcal{S}^{(m)}_{p_{m}}(v+\frac{1}{m})-\mathcal{S}_{p_{m}}^{(m)}(v))$ $=m(\mathcal{S}^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil+1}{m})-\mathcal{S}^{(m)}_{p_{m}}(-\frac{\lceil{mv}\rceil}{m}))$
$=m(G_{\mu\llbracket\partial^{\lceil{mv}\rceil|}p_{m}\rrbracket}(0)-G_{\mu\llbracket\partial^{\lceil{mv}\rceil+1|m}p_{m}\rrbracket}(0))$
$=-G_{\nu^{(m)}_{p_{m}}(v)}(0),$
where $\nu^{(m)}_{p_{m}}(v):=\mu\llbracket\partial^{\lceil{mv}\rceil|}p_{m}\rrbracket-\mu\llbracket\partial^{\lceil{mv}\rceil+1|m}p_{m}\rrbracket$. Observe that this measure has total variation equal to $2d$, it is therefore not uniformly bounded. Nevertheless, we will study the convergence of its Cauchy transform by using the moment method. By definition, the $\ell^{th}$ moment $M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ of $\nu_{p_{m}}^{(m)}(v)$ is
$M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ $=m(M_{\ell}(\partial^{\lceil{mv}\rceil+1|m}p_{m})-M_{\ell}(\partial^{\lceil{mv}\rceil|m}p_{m}))$ (22)
Recall the following asymptotic expansion between the moments $M_{\ell}(p),\,\ell\geq 1$ and the finite free cumulants $\kappa^{(m)}_{\ell}(p),\ell\geq 1$ of a monic polynomial $p$ of degree $m$ (see [arizmendi2018cumulants]):
$M_{\ell}(p)=\sum_{\pi\in{\rm NC}(\ell)}\kappa^{(m)}_{\pi}(p)+\frac{2\ell}{m}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}\kappa^{(m)}_{\pi}(p)+O(\frac{1}{m^{2}})$ (23)
and the following simple formula relating the finite free cumulant of a polynomial $p$ and its derivatives $\partial^{k|m}p$ (see [arizmendi2024s]) :
$\kappa^{(k)}_{\ell}(\partial^{k|m}(p))=(\frac{k}{m})^{\ell-1}\kappa^{(m)}_{\ell}(p),\quad\ell\leq k.$ (24)
Inserting (23) in (22) and using further equation (24), we obtain:
$M_{\ell}(\nu^{(m)}_{p_{m}}(v))$ $=m\bigg(\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}\kappa^{(m)}_{\pi}(p_{m})-(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p_{m})\bigg]$
$\hskip 56.9055pt+2\frac{\ell}{\lceil{mv}\rceil+1}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}\kappa^{(m)}_{\pi}(p_{m})$
$\hskip 142.26378pt-2\frac{\ell}{\lceil{mv}\rceil}\sum_{\begin{subarray}{c}r+s=\ell\\ \pi\in S_{{\rm NC}}(r,s)\end{subarray}}r^{-1}s^{-1}(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p_{m})\bigg)$
$+O(1/m)$
For the first term in the right-hand side of the equation above, with $\kappa_{\pi}(\mu),\pi\in{\rm{\rm NC}}(\ell)$ the partitioned free cumulants of $\mu$:
$m\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil+1}{m})^{\ell-|\pi|}$ $\kappa^{(m)}_{\pi}(p)-(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}\kappa_{\pi}^{(m)}(p)\bigg]$
$=m\sum_{\pi\in{\rm NC}(\ell)}\bigg[(\frac{\lceil{mv}\rceil}{m})^{\ell-|\pi|}((1+\frac{1}{\lceil{mv}\rceil})^{\ell-|\pi|}-1)\kappa^{(m)}_{\pi}(p_{m})\bigg]$
$=\sum_{\begin{subarray}{c}\pi\in{\rm NC}(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O(1/m)$
The sum of the second term and the third term are easily seen to contribute to a factor $O(\frac{1}{m})$. Hence, we obtain for the $\ell^{th}$ moment of $\nu_{p_{m}}^{(m)}(v)$ the following asymptotic expansion:
$M_{\ell}(\nu_{p_{m}}^{(m)}(v))$ $=\sum_{\begin{subarray}{c}\pi\in{\rm NC}(\ell)\\ \pi\neq\hat{0}_{\ell}\end{subarray}}\bigg[(\ell-|\pi|)v^{\ell-|\pi|-1}\kappa_{\pi}(\mu)\bigg]+O({1}/{m})$
$=\partial_{v}\sum_{\pi\in{\rm NC}(\ell)}v^{\ell-|\pi|}\kappa_{\pi}(\mu)+O(1/m)$
$=\partial_{v}m_{\ell}({\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/m)$
$=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})+O(1/m)$
where in the last line $\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}$ is the weak limit of
$\frac{1}{\varepsilon}({\rm Dil}_{v+\varepsilon}\mu^{\boxplus\frac{1}{v+\varepsilon}}-{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})$
as $\varepsilon$ goes to zero. Under our assumptions, weak convergence is equivalent to the convergence of the moments (as an easy application of the Ascoli’s theorem) and thus $\partial_{v}{\rm Dil}_{v}(\mu^{\boxplus 1/v})$ is the unique compactly supported signed measure such that
$M_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})=\partial_{v}M_{\ell}({\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}})$
Let us argue now that the Cauchy transform of $\nu_{p_{m}}^{(m)}(v)$ converges to the Cauchy transform of $\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}$ (analytical in the complementary of an interval $\varepsilon(v),\eta(v)]$ with $\varepsilon(v)>0$ (resp. $\eta(v)$) as a consequence of Lemma 5.6 and the proof of Proposition 5.5 in [arizmendi2024s]; in the sequel $\varepsilon$ (resp. $\eta$) will denote the minimum (resp. maximum) between $\varepsilon(v)$ and $\varepsilon>0$ (resp. between $\eta$ and $\eta(v)$)). We observe first that:
$G_{\nu_{p}(v)}=\{G_{\nu^{(m)}_{p_{m}}(v)},\,m\geq 1\}$
is a normal family of holomorphic functions on the domain $\Omega=[\varepsilon,1]^{c}\subset\hat{\mathbb{C}}$$\mathbb{C}\cup\{$$\infty\}$, by applying Montel’s criterion (Fundamental Normality Test [schiff1993normal]). Let $G\colon\Omega\to\hat{\mathbb{C}}$ be an accumulation point of $G_{\nu_{p}(v)}$. First, $G$ is not identically equal to $\infty$ since $G_{\nu^{(m)}_{p_{m}}(v)}(\infty)=0$ for all $m\geq 1$. Since
$\frac{1}{\ell+1!}\frac{m^{\ell+1}}{dz^{\ell+1}}\biggr|_{z=0}G_{\nu^{(m)}_{p_{m}}(v)}(1/z)=m_{\ell}(\nu_{p_{m}^{(m)}})\rightarrow m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}),$
we infer from the local uniform convergence and the Cauchy formula
$\frac{1}{(\ell+1)!}\frac{m^{\ell}}{dz^{\ell+1}}\biggr|_{z=0}G(1/z)=m_{\ell}(\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}).$
Hence, from unicity of analytic continuation $\{G_{\nu^{(m)}_{p_{m}}(v)},\,m\geq 1\}$ has a unique adherence point. Normality implies that $G_{\nu^{(m)}_{p_{m}}(v)}$ converges toward $G_{\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}$ locally uniformly on $\Omega$. Since $0\in\Omega$, we get, locally uniformly on $v$,
$\nabla^{(m)}\mathcal{S}^{(m)}_{p_{m}}(v)=m(\mathcal{S}^{(m)}_{p_{m}}(v+\frac{1}{m})-\mathcal{S}^{(m)}(v))\rightarrow_{m,+\infty}-G_{\partial_{v}{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(0)=-\partial_{v}G_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(0)=-\partial_{v}{S}_{\mu}(-v).$
The result is proven.
Case 2: The roots of the sequence $(p_{m})_{g\geq 1}$ are contained in $(0,\eta]$.
Let $v\in(0,1-\mu(\{0\}))$. From Theorem 1.2 in [arizmendi2024s] (and the computations done in Case $1$), $\mu\llbracket\partial^{\lceil{mv}\rceil|m}\rrbracket$ tends to ${\rm Dil}_{v}(\mu^{\boxplus\frac{1}{v}})$. Also, by Lemma 5.4 of [arizmendi2024s], for large enough $j$, $\mu\llbracket\partial^{\lceil{mv}\rceil|m}p_{m}\rrbracket$ has support contained in some $\varepsilon,\eta]$. We apply Case $1$ to infer
$\nabla^{(m)}\mathcal{S}^{(m)}_{p_{m}}(-v)=\frac{m}{\lceil{mv}\rceil}\nabla^{(\lceil{mv}\rceil)}\mathcal{S}^{(\lceil{mv}\rceil)}_{\partial^{\lceil{mv}\rceil|m}p_{m}}(-1)=\frac{1}{v}\partial_{z}S_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(-1).$ (25)
Given that $S_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(z)=S_{\mu}(vz)$, we infer
$\frac{1}{v}\partial_{z}S_{{\rm Dil}_{v}\mu^{\boxplus\frac{1}{v}}}(-1)=\partial_{z}S_{\mu}(-v).$
and the result is proven.
Case 3 : Unbounded support
This case is dealt with as in Section 7.4 in [arizmendi2024s].
5.3. Finite free $T$-transform of the Frozen Jacobi process
In this paragraph, we put together the results proved in the last two sections to obtain a differential-difference equation satisfied by the finite free $T$ transform of $\chi^{(r,s,m)}_{t}$ of the Hermitian Jacobi process. Once obtained, we pass to the limit therein and derive a PDE satisfied by the $T$ transform of the free Jacobi process.
Let $e^{(r,s,m)}_{n}(t),0\leq n\leq m,$ be the elementary symmetric functions in the roots of $\chi^{(r,s,m)}(t)$:
$\chi^{(r,s,m)}(t)=\sum_{n=0}^{m}(-1)^{n}e_{n}^{(r,s,m)}(t)x^{m-n},\quad e_{0}^{(r,s,m)}(t)=1.$
The following proposition provides a differential system satisfied by $e^{(r,s,m)}_{n}(t),1\leq n\leq m$. In this respect, the reader readily notices that it is simpler than the one derived in [MR4092353], Lemma 3.1., and corresponding to the Jacobi particle system in [1,1][-1,1].
Proposition 5.3.1.
For any $1\leq n\leq m$ and any time $t>0$:
$\frac{d}{dt}e^{(r,s,m)}_{n}(t)=-n(p+q-(n-1))e_{n}^{(r,s,m)}(t)+(m-(n-1))(p-(n-1))e^{(r,s,m)}_{n-1}(t).$
Proof.
According to Proposition 4.1.1, for any time $t\geq 0$:
$\partial_{t}\chi_{t}^{(r,s,m)}(x)=-m(d-m+1)\chi_{t}^{(r,s,m)}(x)-\mathcal{L}_{x}^{(r,s)}[\chi_{t}^{(r,s,m)}](x).$ (26)
One readily infers from (26):
$\mathcal{L}_{x}^{(r,s)}\chi_{t}^{(r,s,m)}(x)=$ $x(1-x)\partial_{xx}^{2}+[(r+1)-(r+s+2)x]\partial_{x}\chi_{t}^{(r,s,m)}$
$=$ $\sum_{k=0}^{m}(-1)^{k}e^{(r,s,m)}_{k}(t)\left(x(1-x)\partial_{xx}^{2}x^{m-k}+[(r+1)-(r+s+2)x]\partial_{x}x^{m-k}\right)$
$=$ $\sum_{k=0}^{m}(-1)^{k}e^{(r,s,m)}_{k}(t)(m-k)(m-k-1)(x^{m-k-1}-x^{m-k})$
$+\sum_{k=0}^{m}(-1)^{k}e^{(r,s,m)}_{k}(t)(m-k)[(r+1)x^{m-k-1}-(r+s+2)x^{m-k}]$
$=$ $-\sum_{k=0}^{m}(-1)^{k}(m-k)\{(m-k-1)+(r+s+2)\}e^{(r,s,m)}_{k}(t)x^{m-k}$
$+\sum_{k=0}^{m-1}(-1)^{k}(m-k)[r+1+m-k-1]e^{(r,s,m)}_{k}(t)x^{m-k-1}$
$=$ $-\sum_{k=0}^{m}(-1)^{k}(m-k)(m-k+r+s+1)e^{(r,s,m)}_{k}(t)x^{m-k}$
$-\sum_{k=1}^{m}(-1)^{k}(m-k+1)[r+m-k+1]e^{(r,s,m)}_{k-1}(t)x^{m-k}.$
By equating coefficients on both sides of (26), we obtain for any $k\geq 1$:
$\frac{d}{dt}e^{(r,s,m)}_{k}(t)=$ $\left[-m(d-m+1)+(m-k)(m-k+r+s+1)\right]e^{(r,s,m)}_{k}(t)$
$\hskip 142.26378pt+(m-k+1)(r+m-k+1)e^{(r,s,m)}_{k-1}(t)$
$=$ $\left[-m(d-m+1)+(m-k)(p+q-m-k+1)\right]e^{(r,s,m)}_{k}(t)$
$\hskip 142.26378pt+(m-k+1)(p-k+1)e^{(r,s,m)}_{k-1}(t)$
$=$ $-k(p+q-k+1)e^{(r,s,m)}_{k}(t)+(m-k+1)(p-k+1)e^{(r,s,m)}_{k-1}(t).$
Using this proposition, we obtain the following corollary (we recall the notation $\lceil{\cdot}\rceil$ for the ceiling function):
Corollary 5.3.2.
Let $z\in(0,1)$ and $t\geq 0$. Then
$\partial_{t}T_{t}^{(r,s,m)}(z)=[2(m-\lceil{mz}\rceil)-(p+q)]T_{t}^{(r,s,m)}(z)+p-2(m-\lceil{mz}\rceil)\\ +(\frac{\nabla^{(m)}T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})})(p-m+\lceil{mz}\rceil+1))({m-\lceil{mz}\rceil}).$
Proof.
Let $z\in(0,1)$. We put n=m[mz]n=m-[mz] hereafter. We start by differentiating $T^{(r,s,m)}(t)$ in the time variable $t$. Recall:
$T_{t}^{(r,s,m)}(z)=\frac{n}{m-n+1}\frac{e^{(r,s,m)}_{n}(t)}{e^{(r,s,m)}_{n-1}(t)},$
Hence,
$\partial_{t}T^{(r,s,m)}(z)=\frac{n}{(m-n+1)e^{(r,s,m)}_{n-1}(t)}\frac{de^{(r,s,m)}_{n}}{dt}(t)-\frac{T_{t}^{(r,s,m)}(z)}{e^{(r,s,m)}_{n-1}(t)}\frac{de^{(r,s,m)}_{n-1}}{dt}(t).$
Appealing to Proposition 5.3.1, we infer the following equation for the derivative of $T_{t}^{(r,s,m)}(z)$:
$\partial_{t}T_{t}^{(r,s,m)}(z)$ $=(n(n-1)-n(p+q))T_{t}^{(r,s,m)}(z)+{n}(p-(n-1))$
$\hskip 28.45274pt-T_{t}^{(r,s,m)}(z)(n-1)((n-2)-(p+q))-\frac{T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})}(p-(n-2))({n-1}).$
Writing the quotient:
$\frac{T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})}=-\frac{\nabla^{(m)}T_{t}^{(r,s,m)}(z)}{T_{t}^{(r,s,m)}(z+\frac{1}{m})}+1$
and keeping in mind the relation $n-1=m-\lceil{mv}\rceil$, we get the desired result after straightforward computations. ∎
We can now state the main theorem of this section: Theorem 5.3.3 derives the PDE satisfied by the $T$ transform of the free Jacobi process as the limit of the differential-difference equation stated in Proposition 5.3.2.
Theorem 5.3.3.
In the asymptotic regime (2),
  • For each time $t\geq 0$, the $T$ transform $T^{(r,s,m)}_{t/d(m)}(\cdot)$ converges locally uniformly on (1,0)(-1,0).
  • Let $T^{(\lambda,\theta)}\colon\mathbb{R}\times(0,1)$ be the limit, then $T^{(\lambda,\theta)}$ is continuously differentiable and it satisfies:
$\partial_{t}T_{t}^{(\lambda,\theta)}(z)=(2(1-z)\lambda\theta-1)T_{t}^{(\lambda,\theta)}(z)+\theta(1-2\lambda(1-z))$
$\hskip 142.26378pt+\theta(1-\lambda{(1-z)})(1-z){\partial_{z}\log T_{t}^{(\lambda,\theta)}(z)},$
$T_{0}^{(\lambda,\theta)}(z)=1$
Proof.
We already know that for each $t$, the finite free $T$ transform $T^{(r,s,m)}_{t/d(m)}$ converges. By the virtue of Corollary 5.3.2, the only point missing is the local (in time) uniform convergence (in $m$) of $T^{(r,s,m)}_{t/d(m)},~t\geq 0$. We show that the root distribution $\mu^{(r,s,m)}_{t/d(m)}$ converges weakly, locally uniformly in time : if $f:[0,1]\to\mathbb{R}$ is a smooth function, $\mu_{t/d(m)}^{(r,s,m)}(f),\,t\geq 0$ converges locally uniformly in time. The point-wise convergence in time of the $T$-transform implies that for every $t\geq 0$, $\mu_{t/d(m)}^{(r,s,m)}(f)$ converges (in the regime 2). Let us compute the derivative in time:
$\partial_{t}\mu^{(r,s,m)}_{t}(f)=\frac{p}{d(m)}+\frac{p+q}{d(m)}\big\{\frac{1}{m}\sum_{j}x_{j}f^{\prime}(x_{j})\big\}\\ +\frac{1}{2md(m)}\sum_{j\neq k}\frac{f(x_{j})-f(x_{k})}{x_{j}-x_{k}}(x_{j}(1-x_{k})+x_{k}(1-x_{j})).$
Furthermore, we can bound this derivative uniformly on $m$ by
$|\partial_{t}\mu^{(r,s,m)}_{t}(f)|$ $\leq\frac{p}{d(m)}+\frac{p+q}{d(m)}\|f^{\prime}\|_{\infty}+\frac{1}{md(m)}\sum_{k,j}\frac{|f(x_{j})-f(x_{k})|}{|x_{j}-x_{k}|}$
$\leq\frac{p}{d(m)}+\frac{p+q}{d(m)}\|f^{\prime}\|_{\infty}+\frac{m(m-1)}{md(m)}\|f^{\prime\prime}\|_{\infty}.$
We infer that $\{t\mapsto\mu_{t/d(m)}^{(r,s,m)}(f),m\geq 0\}$ is equicontinuous and therefore the local uniform convergence of $\mu^{(r,s,m)}_{t/d(m)}(f)$ as $m$ goes to infinity.
Together we the second Dini’s Theorem (since $T_{t/d(m)}^{(r,s,m)}(\cdot)$ is a non-decreasing step-function), we deduce that the right-hand side of the differential-difference equation of Corollary 5.3.2, if normalize by the factor $d(m)$, converges locally uniformly in time, as follows.
Choosing a $f$ polynomial, we infer that all finite free cumulants of $\chi^{(r,s,m)}_{t}$ converge locally uniformly in time. From the formula recalled in equation (21), we also deduce that all the finite free cumulants of the iterated derivatives of $\chi_{t/d(m)}^{(r,s,m)}$ converge locally uniformly in time (see equations (24)) and therefore the weak convergence, locally uniformly in time, of their root distributions.
The formula (20) will imply local uniform convergence in time of the finite free $S$ and $T$ - transform of $\chi_{t}^{(r,s,m)}$, provided that we can bound away from zero the roots of the iterated derivatives of $\chi_{t}^{(r,s,m)}$, locally uniformly in time. This is done using Lemma 5.5 in [arizmendi2024s], since the cut-off $a$ can be chosen locally uniformly in time.
The limit $T^{(\lambda,\theta)}(z)$ is therefore continuously differentiable in time and for each time $t\geq 0$ and $z\in(-1,0)$:
$\partial_{t}T^{(r,s,m)}_{t/d(m)}(z)\rightarrow T^{(\lambda,\theta)}_{t}(z)$
and we can pass to the limit in both sides of the integro-differential equation of Corollary 5.3.2 to obtain the second assertion of the theorem. ∎